### v1

# MultiSocial: Multilingual Benchmark of Machine-Generated Text Detection of Social-Media Texts

Abstract
Recent LLMs are able to generate high-quality multilingual texts, indistinguishable for humans from authentic human-written ones. Research in machine-generated text detection is however mostly focused on the English language and longer texts, such as news articles, scientific papers or student essays. Social-media texts are usually much shorter and often feature informal language, grammatical errors, or distinct linguistic items (e.g., emoticons, hashtags). There is a gap in studying the ability of existing methods in detection of such texts, reflected also in the lack of existing multilingual benchmark datasets. To fill this gap we propose the first multilingual (22 languages) and multi-platform (5 social media platforms) dataset for benchmarking machine-generated text detection in the social-media domain, called MultiSocial [^1] https://anonymous.4open.science/r/multisocial. It contains 472,097 texts, of which about 58k are human-written and approximately the same amount is generated by each of 7 multilingual LLMs. We use this benchmark to compare existing detection methods in zero-shot as well as fine-tuned form. Our results indicate that the fine-tuned detectors have no problem to be trained on social-media texts and that the platform selection for training matters.
MultiSocial: Multilingual Benchmark of Machine-Generated Text Detection of Social-Media Texts
Dominik Macko, Jakub Kopal, Robert Moro, Ivan Srba Kempelen Institute of Intelligent Technologies, Slovakia {dominik.macko, jakub.kopal, robert.moro, ivan.srba}@kinit.sk
1 Introduction
Figure 1: MultiSocial coverage of languages.
The most advanced text-generation AI models, called large language models (LLMs), are able to generate high-quality texts in various languages (Qin et al., 2024). Although this presents an opportunity to make a human life and work more efficient, it also presents a threat of being misused, as such generated texts are not easily recognisable by humans (Zellers et al., 2019). This is especially crucial in regard to social-media networks (SMN, such as Facebook, X/Twitter, etc.), where anyone can be a source of a harmful content (without editorial consent) with a potentially wide reach (depending on a network) (Aïmeur et al., 2023). To prevent the LLM misuse (e.g., social engineering, disinformation spreading), a reliable mechanism to detect machine-generated text (MGT) is needed.
Unfortunately, the existing research in MGT detection (MGTD) focuses either solely on English (as in most NLP fields, e.g., Dugan et al., 2024) or leaves SMN texts out of scope due to challenges they bring (Kumarage et al., 2024; Lin et al., 2024). These challenges include very informal writing style often used in SMN, such as using slang, ignoring grammar rules, using distinct linguistic items in the texts (e.g., emoticons or hashtags), or usually including very short lengths of the texts.
Our work fills a gap in the state-of-the-art (SOTA) by introducing a new heavily multilingual dataset for MGTD research in the social-media domain. We use this dataset to further benchmark the existing SOTA detectors in various aspects. Specifically, our key contributions include:
(1) The unique multilingual, multi-platform, and multi-generator benchmark dataset of human-written and machine-generated social-media texts, called MultiSocial, covering 22 languages, 5 SMN platforms, and 7 LLM generators.
(2) The first multilingual evaluation of SOTA MGTD methods (statistical, pre-trained, fine-tuned) on social-media texts, focusing on multilingual as well as cross-lingual capability of existing detectors and comparison of different categories. The best detectors perform similarly across all tested languages, although there are still differences between English and non-English in zero-shot evaluation.
(3) The first multi-platform and cross-platform evaluation of SOTA MGTD methods in social-media domain, evaluating differences in MGTD performance based on text types and sources in multiple languages. We found that Telegram offers the best cross-lingual capability.
2 Related Work
Multilingual machine-generated text detection has been gaining attention recently. There have been multiple non-English language MGT detection shared tasks in the last years, such as Russian at RuATD 2022 (Shamardina et al., 2022), Spanish at AuTexTification 2023 (Sarvazyan et al., 2023), Dutch at CLIN33 (Fivez et al., 2024), or multilingual at SemEval-2024 Task 8 (Wang et al., 2024b).
The last one covers 9 languages, and is based on M4GT-Bench (Wang et al., 2024a), multilingual, multi-domain, and multi-generator MGT benchmark. However, its coverage of the SMN domain is rather sparse (solely English-only Reddit texts are included). Moreover, language coverage of the domains is highly imbalanced (e.g., most languages are represented only in the news domain, while Arabic and German also in Wikipedia domain, and Chinese only in the QA domain) and there are only one or two generators used in the multilingual settings. Therefore, cross-lingual evaluation is somewhat limited using such data.
Another benchmark dataset called MULTITuDE (Macko et al., 2023) covers 11 languages; however, 8 of them are included in the test split only with a limited number of samples. Moreover, it is focused on the news domain only, in which the texts are typically long, formal, and carefully checked for grammatical correctness. Such settings are clearly different than those of social media texts. An extension of MULTITuDE dataset to evaluate various authorship obfuscation methods is proposed in (Macko et al., 2024). Although it enables to evaluate robustness of MGT detection methods, it is still limited to news domain.
There is also MAiDE-up dataset (Ignat et al., 2024) of hotel reviews available, covering 10 languages; however, it is limited to GPT-4 generated data only (limiting the generalization of conclusions). Although such texts are more similar to social-media texts than news articles, they cover a single topic (accommodation) and still use a different communication style than the social-media networks. Many other works have focused on detection of fake reviews, but only few of them reflected multilingualism (Duma et al., 2024).
Benchmark datasets HC3 (Guo et al., 2023) and SAID (Cui et al., 2023) contain texts Chinese and English covering forum-like question-answering domain. HC3 contains only ChatGPT-generated machine texts, SAID contains real bot-generated texts (relying on human annotations to identify them). The downside of SAID is that the specific generation model of individual texts is not known. Otherwise, social-media texts are covered only in rather old TweepFake (Fagni et al., 2021) dataset, which includes English-only tweets and cannot be used to evaluate MGT detection methods on social-media texts generated by the most modern LLMs. Fox8-23 (Yang and Menczer, 2023) dataset also focuses on English, and covers presumably only ChatGPT-generated tweets. Similarly, F3 (Lucas et al., 2023) dataset contains English ChatGPT-generated real and fake news as well as tweets. There are other monolingual works, such as (Temnikova et al., 2023) focused on fake tweets in Bulgarian; however, the crafted dataset is not publicly available.
Dataset H/M LLM Lang Domain SMN
TweepFake (Fagni et al., 2021) 12k/12k 6 1 SMN 1
Fox8-23 (Yang and Menczer, 2023) 228k/170k 1 1 SMN 1
F3 (Lucas et al., 2023) 28k/28k 1 1 news, SMN 1
HC3 (Guo et al., 2023) 81k/44k 1 2 QA, Wiki 0
SAID (Cui et al., 2023) 87k/131k N/A 2 QA 0
MULTITuDE (Macko et al., 2023) 8k/66k 8 11 news 0
M4GT-Bench (Wang et al., 2024a) 93k/124k 8 9 6 0
MultiSocial (ours) 58k/414k 7 22 SMN 5
Table 1: Comparison of existing publicly available MGT detection datasets either multilingual or focused on social-media network (SMN) domain. H/M refers to the no. of human-written and machine-generated samples.
Train Test
Language Discord Gab Telegram Twitter WhatsApp all Discord Gab Telegram Twitter WhatsApp all
Arabic (ar) 0 0 7724 7872 0 15596 0 1556 2319 2364 1750 7989
Bulgarian (bg) 0 0 7555 7930 0 15485 0 192 2334 2367 0 4893
Catalan (ca) 6984 0 7157 0 0 14141 2105 1264 2134 1824 144 7471
Chinese (zh) 0 0 7924 0 0 7924 0 830 2380 238 32 3480
Croatian (hr) 7502 0 7065 0 0 14567 2255 1340 2315 91 38 6039
Czech (cs) 3450 0 7690 0 0 11140 2175 492 2309 1017 134 6127
Dutch (nl) 7391 7900 7750 7933 0 30974 2191 2356 2318 2369 397 9631
English (en) 7789 7760 7824 7871 7782 39026 2341 2340 2350 2365 2334 11730
Estonian (et) 6974 0 7520 0 0 14494 2071 805 2259 164 120 5419
German (de) 3407 7863 7880 2095 0 21245 2232 2366 2356 2345 304 9603
Greek (el) 0 0 3814 0 0 3814 0 1195 2274 146 35 3650
Hungarian (hu) 7079 0 7461 0 0 14540 2094 1211 2228 413 22 5968
Irish (ga) 0 0 0 0 0 0 1319 968 821 45 0 3153
Polish (pl) 7158 1829 7733 0 0 16720 2136 2311 2310 172 62 6991
Portuguese (pt) 6860 7916 7842 6481 4354 33453 2284 2371 2347 2360 2363 11725
Romanian (ro) 7436 851 6792 64 0 15143 2236 2349 2298 2378 132 9393
Russian (ru) 0 7875 7827 362 0 16064 0 2355 2340 2361 960 8016
Scottish Gaelic (gd) 0 0 0 0 0 0 150 35 34 0 0 219
Slovak (sk) 0 0 0 0 0 0 107 308 1508 110 0 2033
Slovenian (sl) 0 0 0 0 0 0 203 1912 917 40 0 3072
Spanish (es) 7588 7883 7884 7922 7804 39081 2268 2361 2354 2376 2341 11700
Ukrainian (uk) 0 0 7802 0 0 7802 0 174 2342 70 0 2586
Total 79618 49877 133244 48530 19940 331209 28167 31091 44847 25615 11168 140888
Table 2: MultiSocial text sample counts across languages and platforms for train and test split.
Table 1 includes comparison of the new dataset proposed in this work (described in the next section) with the selected existing publicly available datasets for MGT detection.
3 Dataset
Since there is no dataset of multilingual SMN texts containing human-written texts along with the texts generated by SOTA text-generation LLMs, we have crafted a new MultiSocial benchmark dataset. It contains human-written data from five different social-media platforms (reused from the existing multilingual SMN datasets), namely Telegram, Twitter (X), Gab, Discord, and WhatsApp, including post-like as well as chat-like texts. For each authentic human-written text, the texts generated by 7 SOTA LLMs are included by using three iterations of paraphrasing. Details regarding dataset construction (including selection, pre-processing, generation, and post-processing of texts) are provided in Appendix B.
Dataset consists of 472,097 texts (58k are human-written) split into train and test subsets, of which sample counts are summarized across languages and platforms in Table 2. We have conducted linguistic analysis of the machine-generated texts along with the similarity comparison to human texts, with a manual human check of a balanced subset, and identified small portion (about 1%) of noise in the generated data (e.g., “As an AI model…”), indicating model failure during generation. We have intentionally left such samples in the dataset (clearly marked in the data) for further analysis purpose (as indicated in Appendix B). We however filter-out the identified noise for the purpose of the experiments in this study. Although such a post-processing cannot guarantee 100% removal of noisy data, the obvious failures are cleaned.
Language Selection.
The MultiSocial benchmark dataset covers 22 (high- and low-resource) languages (some only in the testing split), 20 of which (18 of Indo-European and 2 of Uralic language families) have been selected based on our research-projects needs. However, we have intentionally included 2 more (Arabic of Semitic and Chinese of Sino-Tibetan family), which are completely linguistically and geographically unrelated, to study cross-lingual characteristics (Figure 1).
The dataset is strongly focused on the Indo-European language family, but contains 4 language families in total. Test split includes all of the train languages, but also 2 Celtic languages, 1 more South Slavic and 1 more West Slavic language. There are 5 writing scripts in both train and test splits of the dataset, where majority of languages uses Latin, but there is also Cyrillic (Russian, Ukrainian, Bulgarian), Arabic, Hanzi (Chinese), and Greek. Nice feature is that the train split includes at least 3 representatives of Germanic, Romance, Slavic-Latin, and Slavic-Cyrillic, which enables various combinations of studies regarding multilingual and cross-lingual characteristics of machine-generated text detection. Furthermore, in Slavic and Romance languages, there are included at least 2 representatives of languages that can be considered high-resource and low-resource.
Although not all languages have enough samples from each of five platforms, specific subsets can be used to study specific characteristics (e.g., cross-platform transferability using English and Spanish languages, cross-lingual transferability using Telegram platform, surprise language and/or surprise platform evaluation not using specific portions of the train data, etc.).
4 Detection Methods
For the benchmark purpose, we have covered 3 specific categories of detection methods: statistical zero-shot (methods using statistical differences to differentiate human-written and machine-generated texts applicable without training), pre-trained (directly applicable models that are fine-tuned for the MGT detection task using different data – i.e., out-of-distribution), and fine-tuned (foundation models fine-tuned for the MGT detection task using MultiSocial dataset – i.e., in-distribution).
For statistical category, we have selected the following 5 most-promising methods (excluding perturbation-based and multi-generation methods due to high computing costs): Binoculars (Hans et al., 2024) with Falcon-7B (Almazrouei et al., 2023) as an observer model and Falcon-7B-Instruct as a performer model, Fast-DetectGPT (Bao et al., 2023) with GPT-J-6B (Wang and Komatsuzaki, 2021) as both the reference and sampling models, LLM-Deviation (Wu and Xiang, 2023), DetectLLM-LRR (Su et al., 2023), S5 (Spiegel and Macko, 2024) (multiplying 5 statistical metrics of Likelihood, Entropy, Rank, Log-Rank, and LLM-Deviation), all three of them using GPT-J-6B as a base model (the same as in Fast-DetectGPT).
For pre-trained category, we have selected the following 5 detectors with a multilingual potential: ChatGPT-detector-RoBERTa-Chinese (Guo et al., 2023) (with a Chinese fine-tuned model), Longformer Detector (Li et al., 2023) (showing decent multilingual potential in Macko et al., 2024), ruRoBERTa-ruatd-binary [^2] https://huggingface.co/orzhan/ruroberta-ruatd-binary (as a best single-model system in RuATD 2022 Shamardina et al., 2022), BLOOMZ-3B-mixed-detector (Nicolai Thorer Sivesind and Andreas Bentzen Winje, 2023) (with a heavily multilingual fine-tuned LLM), and RoBERTa-Large OpenAI Detectors (Solaiman et al., 2019) (as a widely used representative, although English-only).
For fine-tuned category, we have selected 7 multilingual foundational models covering multiple architectures and model sizes: mDeBERTa-v3-base (He et al., 2022) (as the best detector in Macko et al., 2023), XLM-RoBERTa-large (Conneau et al., 2020) (as the best detector in Macko et al., 2023 based on AUC ROC), Mistral-7B (Jiang et al., 2023) (as the best single-model multilingual detector in of SemEval-2024 Task 8 Wang et al., 2024b), Llama-3-8B (AI@Meta, 2024) (as a SOTA multilingual smaller LLM), Aya-101 (Üstün et al., 2024) (as a SOTA representative of multilingual LLMs with encoder-decoder architecture), BLOOMZ-3B (Muennighoff et al., 2022) (as the best pre-trained detector base model), and Falcon-rw-1B (as a smaller version of the best performing model at ALTA 2023 Gagiano and Tian, 2023, since the 7B model version is already covered by better performing Mistral, Spiegel and Macko, 2024).
5 Experimental Results
Rank Detector AUC ROC
MacroF1
@5%FPR
1 Llama-3-8b-MultiSocial 0.9769 0.8696
2 Mistral-7b-v0.1-MultiSocial 0.9768 0.8692
3 Aya-101-MultiSocial 0.9731 0.8462
4 Falcon-rw-1b-MultiSocial 0.9592 0.7810
5 BLOOMZ-3b-MultiSocial 0.9582 0.7843
6 XLM-RoBERTa-large-MultiSocial 0.9553 0.7840
7 mDeBERTa-v3-base-MultiSocial 0.9544 0.7652
8 BLOOMZ-3b-mixed-Detector 0.7553 0.3024
9 DetectLLM-LRR 0.7464 0.2523
10 LLM-Deviation 0.7454 0.2497
11 S5 0.7418 0.2465
12 Fast-Detect-GPT 0.7418 0.3605
13 Binoculars 0.7248 0.2815
14 ChatGPT-Detector-RoBERTa-Chinese 0.7180 0.3416
15 ruRoBERTa-ruatd-binary 0.4817 0.1711
16 Longformer Detector 0.4615 0.1516
17 RoBERTa-large-OpenAI-Detector 0.3450 0.1376
Table 3: Benchmark of the selected MGTD methods of statistical, pre-trained, and fine-tuned categories.
Test Language [AUC ROC]
Category Detector ar bg ca cs de el en es et ga gd hr hu nl pl pt ro ru sk sl uk zh all
P BLOOMZ-3b-mixed-Detector 0.79 0.74 0.79 0.80 0.77 0.76 0.80 0.79 0.83 0.78 0.66 0.77 0.84 0.75 0.78 0.79 0.66 0.68 0.76 0.64 0.70 0.69 0.76
ChatGPT-Detector-RoBERTa-Chinese 0.72 0.80 0.76 0.66 0.80 0.75 0.90 0.81 0.82 0.73 0.63 0.63 0.78 0.70 0.62 0.73 0.75 0.76 0.63 0.66 0.74 0.81 0.72
Longformer Detector 0.34 0.48 0.32 0.47 0.43 0.54 0.65 0.43 0.48 0.53 0.49 0.51 0.50 0.41 0.46 0.42 0.41 0.46 0.45 0.46 0.61 0.47 0.46
RoBERTa-large-OpenAI-Detector 0.73 0.43 0.43 0.14 0.32 0.74 0.52 0.30 0.20 0.30 0.48 0.19 0.13 0.30 0.21 0.23 0.24 0.54 0.26 0.33 0.36 0.60 0.35
ruRoBERTa-ruatd-binary 0.40 0.63 0.56 0.43 0.43 0.35 0.56 0.47 0.43 0.49 0.47 0.48 0.46 0.50 0.44 0.43 0.34 0.70 0.45 0.47 0.59 0.44 0.48
S Binoculars 0.70 0.68 0.62 0.74 0.75 0.79 0.80 0.76 0.74 0.71 0.71 0.79 0.78 0.72 0.75 0.75 0.74 0.72 0.71 0.73 0.64 0.74 0.72
DetectLLM-LRR 0.79 0.86 0.69 0.93 0.78 0.88 0.80 0.82 0.88 0.79 0.74 0.88 0.94 0.79 0.88 0.84 0.87 0.78 0.85 0.79 0.75 0.78 0.75
Fast-Detect-GPT 0.75 0.65 0.61 0.81 0.77 0.66 0.80 0.74 0.69 0.70 0.74 0.80 0.77 0.74 0.77 0.77 0.77 0.73 0.71 0.74 0.70 0.74 0.74
LLM-Deviation 0.82 0.86 0.68 0.93 0.79 0.89 0.80 0.82 0.90 0.81 0.79 0.89 0.94 0.79 0.89 0.84 0.88 0.78 0.86 0.81 0.75 0.79 0.75
S5 0.80 0.85 0.68 0.92 0.78 0.88 0.77 0.81 0.89 0.80 0.78 0.88 0.94 0.77 0.88 0.83 0.88 0.78 0.85 0.80 0.74 0.78 0.74
Table 4: Per-language AUC ROC performance of zero-shot statistical (S) and pre-trained (P) MGT detectors. The data are too difficult for the three under-performing pre-trained models.
Test Language [AUC ROC mean]
Category Platform ar bg ca cs de el en es et ga gd hr hu nl pl pt ro ru sk sl uk zh all
P Discord N/A N/A 0.92 0.81 0.86 N/A 0.91 0.89 0.86 N/A N/A 0.75 0.84 0.81 0.81 0.82 0.81 N/A N/A N/A N/A N/A 0.82
Gab N/A N/A N/A N/A 0.71 N/A 0.82 0.71 N/A N/A N/A N/A N/A 0.71 0.64 0.72 0.61 0.62 N/A N/A N/A N/A 0.66
Telegram 0.73 0.77 0.73 0.73 0.79 0.81 0.86 0.80 0.84 N/A N/A 0.71 0.86 0.68 0.70 0.77 0.77 0.76 N/A N/A 0.73 0.76 0.74
Twitter 0.84 0.79 N/A N/A 0.78 N/A 0.85 0.73 N/A N/A N/A N/A N/A 0.73 N/A 0.75 0.65 0.80 N/A N/A N/A N/A 0.73
WhatsApp N/A N/A N/A N/A N/A N/A 0.82 0.86 N/A N/A N/A N/A N/A N/A N/A 0.76 N/A N/A N/A N/A N/A N/A 0.79
all 0.76 0.77 0.77 0.73 0.78 0.75 0.85 0.80 0.82 0.75 N/A 0.70 0.81 0.72 0.70 0.76 0.71 0.72 0.70 0.65 0.72 0.75 0.74
S Discord N/A N/A 0.85 0.90 0.87 N/A 0.90 0.89 0.86 N/A N/A 0.89 0.91 0.89 0.92 0.90 0.89 N/A N/A N/A N/A N/A 0.88
Gab N/A N/A N/A N/A 0.73 N/A 0.76 0.74 N/A N/A N/A N/A N/A 0.72 0.77 0.73 0.74 0.71 N/A N/A N/A N/A 0.69
Telegram 0.76 0.78 0.62 0.88 0.71 0.86 0.81 0.77 0.84 N/A N/A 0.89 0.91 0.74 0.84 0.82 0.87 0.74 N/A N/A 0.71 0.75 0.74
Twitter 0.76 0.79 N/A N/A 0.87 N/A 0.83 0.72 N/A N/A N/A N/A N/A 0.78 N/A 0.86 0.85 0.90 N/A N/A N/A N/A 0.74
WhatsApp N/A N/A N/A N/A N/A N/A 0.69 0.85 N/A N/A N/A N/A N/A N/A N/A 0.78 N/A N/A N/A N/A N/A N/A 0.72
all 0.77 0.78 0.65 0.87 0.77 0.82 0.79 0.79 0.82 0.76 N/A 0.85 0.87 0.76 0.83 0.81 0.83 0.76 0.80 0.77 0.71 0.76 0.74
Table 5: Per-platform mean AUC ROC performance of well-performing zero-shot MGT detectors per category. N/A refers to not enough samples (at least 2000) in MultiSocial for a combination of language and platform. Discord data are the easiest for the detection, Gab data are the most difficult.
Firstly, we provide benchmark evaluation of the selected MGTD methods on all MultiSocial test data (Table 3). It provides a fair comparison of the methods, although the data samples among platforms and languages are not perfectly balanced. Therefore, for further experiments targeting specific cross-lingual and cross-platform research questions, the carefully selected parts of train and test splits are used (described in the corresponding subsections). Per-language, per-platform, and per-generator results are provided in Appendix F. Although worse-than-random performances of some pre-trained detectors indicate a potential problem with the detection (e.g., flipped labels), it is not the case as confirmed by the results in Table 17.
For comparison of MGTD methods, we use AUC ROC (area under the curve of receiver operating characteristic) as a classification-threshold independent metric (not affected by a threshold calibration on in-domain data), also used by (Hans et al., 2024). Due to imbalanced test data (machine class contains 7x more samples), we also use Macro avg. F1-score @ 5% FPR (false positive rate) as a metric balancing between a precision and a recall of the classification, while the threshold is calibrated based on the train data (to avoid data leakage) ROC curve to achieve 5% FPR (similarly used in Dugan et al., 2024).
Since Gemini-generated data have used a slightly different generation process (e.g., jailbreak prompt, see Appendix B) and achieved highly outlier word-count and unique-words scores in Table 11, we do not use Gemini-generated data in the training (fine-tuning or classification-threshold calibration). Nevertheless, they are still included in the evaluation (can serve for unseen generator evaluation).
5.1 Multilingual Zero-shot Detection
This experiment is focused on the following research question: RQ1: How well are social-media texts of multiple languages and platforms detectable by MGT detectors applicable in zero-shot manner (out-of-distribution, without further in-domain training)? Since SMN texts are usually shorter than commonly used news articles, the detection performance of existing directly usable (i.e., without in-domain training) MGT detectors is still unknown (it could differ from the reported performance on other domains). This has not been evaluated in the multilingual settings. Are there differences in multilingual MGT detection among different sources of SMN content (e.g., Twitter vs. Telegram vs. Gab)? Is there a difference between statistical (could be language independent) and pre-trained (heavily dependent on pre-training languages) zero-shot detectors?
To answer these questions, we compare the per-language performance of statistical and pre-trained MGTD methods (collectively called zero-shot methods for this purpose) based on AUC ROC (to avoid in-domain touch with the data) in Table 4. To compare per-language performance per platforms, we consider only cases where there are at least 2000 samples available per platform and language (approximately 250 texts per each generator). The summarized results of the comparison are provided in Table 5. The all row represents performance of MGT detectors of the corresponding category for all platforms data combined (excluding only results for Scottish due to not having enough samples). Due to low performance of three out of five selected pre-trained detectors (see Table 3), we average results for this category only for the two well-performing detectors (BLOOMZ-3b-mixed-Detector and ChatGPT-Detector-RoBERTa-Chinese). Full results are available in Appendix F.
Test Language [AUC ROC]
Detector ar bg ca cs de el en es et $\star$ $\star$ hr hu nl pl pt ro ru $\star$ $\star$ uk zh all
Aya-101-MultiSocial 0.97 0.99 0.97 0.98 0.97 0.97 0.98 0.98 0.98 0.95 0.92 0.98 0.99 0.97 0.98 0.98 0.98 0.96 0.98 0.95 0.95 0.97 0.97
BLOOMZ-3b-MultiSocial 0.96 0.98 0.96 0.97 0.95 0.96 0.98 0.97 0.98 0.90 0.82 0.96 0.99 0.94 0.95 0.97 0.95 0.94 0.95 0.88 0.90 0.97 0.96
Falcon-rw-1b-MultiSocial 0.95 0.98 0.97 0.97 0.96 0.96 0.98 0.96 0.98 0.92 0.87 0.96 0.99 0.95 0.96 0.96 0.95 0.94 0.95 0.87 0.91 0.96 0.96
Llama-3-8b-MultiSocial 0.97 0.99 0.98 0.99 0.98 0.97 0.99 0.98 0.99 0.94 0.90 0.98 0.99 0.98 0.98 0.98 0.98 0.96 0.98 0.95 0.95 0.98 0.98
Mistral-7b-v0.1-MultiSocial 0.97 0.99 0.98 0.99 0.98 0.97 0.99 0.98 0.98 0.93 0.93 0.99 1.00 0.97 0.98 0.98 0.97 0.97 0.97 0.94 0.96 0.98 0.98
XLM-RoBERTa-large-MultiSocial 0.95 0.98 0.94 0.98 0.96 0.95 0.96 0.96 0.97 0.88 0.78 0.97 0.99 0.95 0.97 0.95 0.96 0.95 0.96 0.91 0.92 0.93 0.96
mDeBERTa-v3-base-MultiSocial 0.94 0.98 0.94 0.97 0.95 0.94 0.96 0.96 0.98 0.90 0.79 0.97 0.99 0.95 0.96 0.96 0.96 0.93 0.96 0.92 0.93 0.94 0.95
Table 6: Per-language AUC ROC performance of fine-tuned MGT detectors. $\star$ marks languages not in train set. Larger models achieve better performance.
To evaluate whether the differences between mean AUC ROC of statistical and the best pre-trained MGT detectors are statistically significant, we conduct paired t-tests for each test language and check whether p-value is < 0.05. Analogously, we have verified significance of differences between per-platform means in each category.
There are differences in performance of SOTA zero-shot MGTD methods on texts of English and non-English languages. When considering the well-performing pre-trained and all statistical detectors, the difference between performances on English and combined non-English texts is statistically significant (higher on English). However, Longformer Detector clearly performed better on English than the other languages. Similarly, ruRoBERTa performed better on Russian, Ukrainian, and Bulgarian than the others. OpenAI Detector has clearly not been trained on SMN texts, since not performing well even in English (there are also huge differences in performance based on the generators, see Table 17). The other two pre-trained and all statistical detectors performed similarly across languages, although the Chinese detector performed worse on Slavic-Latin languages.
There are significant differences in performance of SOTA zero-shot MGTD methods on texts of different platforms. The detectors are able to better detect MGT of Discord SMN than the others (although the differences to other platforms are not statistically significant for pre-trained detectors). On the other hand, the Gab texts are the most difficult for them to classify (although the differences to Telegram for pre-trained and WhatsApp for statistical detectors are not statistically significant). There is no clear indication for the length of such texts affecting these results, since Discord has the lowest (9) and WhatsApp and Twitter the highest (18) median value of word-count text length. We can speculate that since Gab is known to have more toxic content (vulgarisms, hate speech), it can be more difficult for detection.
There are negligible differences in performance of SOTA zero-shot statistical and best pre-trained MGTD methods. When considering the two best performing pre-trained detectors, which achieved 0.72-0.76 AUC ROC (0.62-0.9 in per-language evaluation), the performance is competitive with the statistical detectors, achieving 0.72-0.75 AUC ROC (0.61-0.94 in per-language evaluation). In regard to multilingual performance, the statistical detectors tends to achieve higher performance for Slavic-Latin and Uralic languages (confirmed by Telegram-only data), under-performing for Catalan. This is not the case of pre-trained detectors, under-performing for Scottish and Slovenian, and the Chinese detector shows rather opposite patterns for Slavic-Latin languages. The t-tests confirmed that the differences between these two categories are statistically significant only for Catalan, Scottish and Slovenian languages.
Train Test Language [AUC ROC mean (±confidence interval)]
Language ar bg ca cs de el en es et ga gd hr hu nl pl pt ro ru sk sl uk zh all
en 0.81 (±0.05) 0.90 (±0.08) 0.78 (±0.03) 0.91 (±0.05) 0.88 (±0.02) 0.90 (±0.03) 0.96 (±0.01) 0.87 (±0.06) 0.92 (±0.03) N/A N/A 0.94 (±0.03) 0.97 (±0.03) 0.84 (±0.02) 0.89 (±0.05) 0.92 (±0.02) 0.94 (±0.02) 0.87 (±0.05) N/A N/A 0.81 (±0.05) 0.73 (±0.14) 0.87 (±0.04)
es 0.82 (±0.05) 0.89 (±0.08) 0.85 (±0.03) 0.89 (±0.06) 0.89 (±0.03) 0.87 (±0.06) 0.90 (±0.04) 0.94 (±0.01) 0.90 (±0.05) N/A N/A 0.92 (±0.05) 0.95 (±0.04) 0.84 (±0.03) 0.88 (±0.06) 0.93 (±0.02) 0.93 (±0.03) 0.88 (±0.04) N/A N/A 0.82 (±0.05) 0.73 (±0.14) 0.86 (±0.05)
ru 0.81 (±0.10) 0.93 (±0.05) 0.76 (±0.10) 0.87 (±0.08) 0.84 (±0.04) 0.88 (±0.06) 0.87 (±0.06) 0.82 (±0.11) 0.88 (±0.08) N/A N/A 0.89 (±0.07) 0.91 (±0.07) 0.79 (±0.06) 0.87 (±0.07) 0.86 (±0.07) 0.88 (±0.07) 0.94 (±0.02) N/A N/A 0.89 (±0.04) 0.73 (±0.17) 0.84 (±0.08)
en-es-ru 0.89 (±0.03) 0.94 (±0.04) 0.86 (±0.03) 0.93 (±0.03) 0.91 (±0.03) 0.93 (±0.02) 0.95 (±0.01) 0.93 (±0.02) 0.93 (±0.03) N/A N/A 0.94 (±0.04) 0.97 (±0.02) 0.86 (±0.03) 0.91 (±0.04) 0.94 (±0.01) 0.95 (±0.02) 0.93 (±0.03) N/A N/A 0.89 (±0.03) 0.86 (±0.08) 0.91 (±0.03)
Table 7: Cross-lingual mean AUC ROC performance of the selected MGT detectors fine-tuned monolingually (en, es and ru) and multilingually (en-es-ru), evaluated based on Telegram data (for training as well as for testing), reported along with 95% confidence interval error bounds. N/A refers to not enough samples (at least 2000) in MultiSocial Telegram data. Multilingual fine-tuning helps especially for languages unrelated to train languages.
5.2 Multilingual Fine-tuned Detection
This experiment is focused on the following research question: RQ2: How well are social-media texts of multiple languages detectable by fine-tuned MGT detectors? SMN texts have a higher variety of styles and lower grammatical correctness than news articles. Are language models able to be fine-tuned for the MGT detection task using such texts? Also, it is unknown which detection method is the most universal in regard to the diversity of use cases (different text lengths, different sources). Is the best MGT detector for news articles the same as for SMN content? Is the transferability to different languages the same?
Similarly to the previous experiment, we firstly compare AUC ROC performance per each test language in Table 6. The foundational models are fine-tuned in this experiment using all MultiSocial train data (except the samples generated by Gemini). The results show only small differences among the selected detectors, with pretty much steady performance across languages. When looking at the per-generator performance in Table 18, we might observe slightly decreased performance for Gemini (as not used for training), but also for OPT-IML-Max-30b and Aya-101, both having a shorter word count text lengths (Table 11).
The multilingual models are able to be fine-tuned for MGTD task in social-media domain. The performance reached above 0.9 AUC ROC in all train languages, with slightly lower performance of some models in test-only languages (Scottish and Slovenian). Therefore, the style and form of the SMN texts does not seem to limit the ability of the models to serve as fine-tuned detectors.
For the cross-lingual evaluation, we use the same language setting as used by MULTITuDE (Macko et al., 2023), which was focused on news domain, for the results to be comparable. Specifically, we use English, Spanish and Russian Telegram data (having enough samples for training, approximately the same size, representatives of different language-family branches) for monolingual as well as multilingual fine-tuning (per-language pseudo-random sub-sampling to 1/3 of the samples count to reach the same cumulative count as in monolingual fine-tuning). Due to lower number of samples in the selected portions of the train dataset than in using full data in the previous experiment, we prolong the fine-tuning process to 7 epochs for models to be able to train well. The cross-lingual results are summarized in Table 7, where mean performances across detectors are reported (per-detector results are provided in Appendix F). As the per-detector results in Table 25 clearly indicate different behaviour of some detectors across languages, we provide an ablation study in Appendix E, where we aggregate the results per the two identified groups of detectors.
Multilingual fine-tuning can improve cross-lingual transferability. Our experiments show that fine-tuning using multiple languages is almost always superior to monolingual setting (see Table 7). However, the rate of improvement can vary depending on model architecture and train-test language similarity. We can observe more noticeable improvements on unrelated languages such as Arabic and Chinese. The ablation study revealed that there is a subset of detectors (with non-autoregressive models) for which the differences between monolingually and multilingually fine-tuned versions are not statistically significant for any language.
Train Test Platform [AUC ROC mean (±confidence interval)]
Platform Discord Gab Telegram Twitter WhatsApp    all
Discord 0.98 (±0.00) 0.84 (±0.02) 0.88 (±0.02) 0.82 (±0.04) 0.89 (±0.03) 0.88 (±0.02)
Gab 0.96 (±0.01) 0.94 (±0.01) 0.93 (±0.02) 0.94 (±0.03) 0.91 (±0.02) 0.93 (±0.01)
Telegram 0.98 (±0.00) 0.92 (±0.02) 0.96 (±0.01) 0.95 (±0.02) 0.95 (±0.01) 0.95 (±0.01)
Twitter 0.97 (±0.01) 0.91 (±0.01) 0.92 (±0.02) 0.98 (±0.01) 0.92 (±0.02) 0.93 (±0.01)
WhatsApp 0.97 (±0.01) 0.90 (±0.01) 0.93 (±0.01) 0.92 (±0.02) 0.97 (±0.01) 0.93 (±0.01)
all 0.98 (±0.01) 0.93 (±0.02) 0.95 (±0.01) 0.96 (±0.01) 0.95 (±0.01) 0.95 (±0.01)
Table 8: Cross-platform mean AUC ROC performance of the selected fine-tuned MGT detectors, reported along with 95% confidence interval error bounds. Telegram-based mono-platform fine-tuning shows the best cross-platform transferability.
For the cross-platform evaluation, we use English and Spanish combined data only, since they are balanced across all the platforms and have enough samples for training. We fine-tuned the selected models in mono-platform (a single SMN platform data) and multi-platform (all platforms combined; similarly to the previous expriment, we have used per-platform pseudo-random sub-sampling to 1/5 of the samples count to reach the same cumulative count as in mono-platform fine-tuning) manner. The per-test-platform results are summarized in Table 8, where mean performances across detectors are reported for each train platform (per-detector results are provided in Appendix F). Gab platform data are still the most difficult for MGT detection and Discord data are the easiest.
Using Telegram data for mono-platform fine-tuning achieves the best cross-platform transferability of detection performance. We can speculate that the reason may be the well-diversified texts across lengths and forms. On the other hand, using Discord data achieves the worst cross-platform transferability. Similarly to zero-shot detectors, there are significant differences among different platforms data disregarding mono-platform and multi-platform fine-tuning. On average, the multi-platform fine-tuning could not reach the performance of mono-platform fine-tuning in the in-platform evaluation. Although, beside the Telegram-trained detectors, the multi-platform ones reached the best performance across platforms. The differences between these two versions (Telegram and all) are not statistically significant for any test platform.
6 Discussion
Shorter and more informal style of the texts in social-media domain does not prevent detectors to be fine-tuned for this domain with superior performance. Despite our assumptions of SMN texts to be quite challenging for fine-tuned detectors, the results indicate that they have no problem to be trained on such texts. The best fine-tuned detectors achieved 0.98 AUC ROC performance and 0.87 Macro average F1-score, with a steady performance across all the test languages.
Bigger LLMs achieve higher performance as fine-tuned MGT detectors than smaller foundational models. The size of the models seems to affect their performance, as >7b parameters models achieved significantly superior performance compared to <7b models. However, for practical application, one must consider a trade-off between detection performance and inference costs, since even the smallest mDeBERTa-v3-base achieved much better performance than zero-shot detectors (which also use base models of >6b parameters).
Multilingual fine-tuning helps cross-lingual transferability of autoregressive models in the MGTD task. We have noticed a clear difference in performances of two groups of fine-tuned detectors, namely the foundational models with autoregressive pre-training and the models with autoencoding (for masked language modeling, XLM-RoBERTa and mDeBERTa) or sequence-to-sequence (Aya) pre-training. The ablation study (Appendix E) aggregating the results for these two groups revealed that the benefit of multilingual fine-tuning is higher in autoregressive group than the other, where the differences are not statistically significant. Also, linguistical similarity between languages seems to affect the transferability in the autoregressive group more intensively.
Selection of social-media platform for fine-tuning matters. There are also significant differences between models trained on single vs multiple platform dataset. For example, on Twitter data, the Discord-trained detectors achieve on average 13% lower AUC ROC than Telegram-trained detectors.
The best detectors fine-tuned on social-media texts still outperform zero-shot detectors on news-domain texts generated by the same models. Although there is a drop in such out-of-domain performance (Appendix D), the detection ability in most languages is still better than that of zero-shot detectors if the data are generated by the generators used for training (i.e., cross-domain). If a different generator is used (i.e., cross-domain and cross-generator), the Fast-Detect-GPT and Binoculars can outperform the fine-tuned detectors.
7 Conclusions
We have created and published a unique multi-platform and massively multilingual dataset, named MultiSocial, to benchmark machine-generated text detection methods on social-media texts. It covers 7 most modern text-generation AI models (of various sizes and architectures), 5 social-media network platforms, and 22 languages of 4 primary language families. We have used this dataset to benchmark 17 carefully selected state-of-the-art detection methods of 3 categories (statistical zero-shot, pre-trained, and fine-tuned) and compare their multi-platform and multi-lingual capabilities (as well as cross-lingual and cross-platform capabilities of fine-tuned detectors). We have discussed the most interesting findings, including that the detection models can be fine-tuned to the machine-generated text detection task using social-media texts (shorter lengths, informal style, emoticons and hashtags) quite well, with the performance comparable to the performance reported in other domains (e.g., news articles).
Due to rather high performance differences in cross-platform evaluation, the further work should be focused to a more-detailed analysis of cross-domain multilingual capability of the state-of-the-art detectors. The proposed MultiSocial dataset can be used for a more detailed multilingual evaluation as well, such as a selection of optimal minimal subset of languages and platforms for training.
Limitations
Limited text generation models and approaches. We have used 7 SOTA LLMs of various architectures and sizes for the text generation. However, these can not cover the huge amount and variety of different text-generation models available (with new models coming each month). We have selected the 3-iteration paraphrasing approach for the text generation. There are other approaches usable for the generation of social-media texts (we have experimented with few of them) that could yield different results of the benchmark.
Limited selection of machine-generated text detection methods. We have selected 17 detectors for the benchmark evaluation. There exist other MGT detection methods (e.g., perturbation or multi-generation based statistical methods or non-zero-shot statistical methods) that have not been included due to cost-efficiency of their usage. We have also not included combinations of multiple methods into the benchmark comparison.
Limited scope of the experiments. Given the multipurpose nature of the proposed MultiSocial benchmark dataset, there are plenty of other research questions that can be targeted, such as the most effective minimal combination of train languages to reach a certain cross-lingual capability. Since we are publishing the MultiSocial dataset as well as the code used in our benchmark, our results are fully reproducible and further research questions can be easily targeted by fellow researchers and future works.
Ethics Statement
Intended Use. We have proposed a MultiSocial dataset along with a code for benchmark of multilingual machine-generated text detection methods. The released artefacts are intended for research purpose only. They are not intended for deployment of actual services making automated decisions, as the classifications are not fully reliable, and could potentially do harm (e.g., false positive prediction, where human-written text is classified as machine generated).
Failure Modes. As confirmed by our experiments, although the detectors can generalize to data from other platforms, languages, generators or domains, this capability is limited and we do observe differences. The behavior on data from other platforms, languages, etc. is thus unknown and should be properly tested before any use.
Biases. Although the dataset contains a wide variety of languages (22 in total) covering various scripts and language families (see Section 3), the dataset is still biased towards Indo-European languages with 18 out of 22 belonging to this family. The dataset also reflects the topics characteristic for the time and the social media included in the 6 original datasets used as sources of human-written texts, but they should already be rather varied due to sheer volume of the data included (see Appendix B).
Misuse Potential. We work with already publicly available datasets of human-written social media content as well as with publicly available LLMs to generate the texts. In general, the human-written texts are not specifically targeted on disinformation, sensitive or toxic content, but the presence of such content (especially the toxic content in case of Gab) cannot be ruled out. Secondly, although we have revealed in the paper which languages are more difficult for the SOTA detection methods to be applied in (i.e., a misuse potential of LLM-generated texts in those languages is higher), the overall misuse potential of our work is rather limited. To the opposite, our work aims to increase the robustness and generalizability of the current SOTA detection methods.
Collecting Data from Users. We have not collected any user data as a part of this work, but are re-using already publicly available datasets of social media posts.
Potential Harm to Vulnerable Populations. We are not aware of any potential harms unless the detectors were employed outside of their intended use, where they could potentially flag also legitimate uses of machine-generated text.
Licensing. As already mentioned, MultiSocial dataset is based on human data of 6 existing datasets. We have made sure to use and re-publish the data in accordance with their licenses. Specifically, two of the datasets are licensed by CC BY 4.0, one by AGPL-3.0, two for research purpose only, and one with no explicit licensing (thus assumed copyrighted). All of such licensing allows use of data for non-commercial research such as our work. We have also checked and followed licensing and terms of use of the used text generation LLMs. Therefore, we release the MultiSocial data with attribution to the sources of human texts for non-commercial research purpose only.
Acknowledgments
This work was partially supported by the projects funded by the European Union under the Horizon Europe: AI-CODE, GA No. 101135437, VIGILANT, GA No. 101073921; and by Modermed, a project funded by the Slovak Research and Development Agency, GA No. APVV-22-0414.
Part of the research results was obtained using the computational resources procured in the national project National competence centre for high performance computing (project code: 311070AKF2) funded by European Regional Development Fund, EU Structural Funds Informatization of Society, Operational Program Integrated Infrastructure.
Appendix A Computational Resources
For social-media texts generation and similarity-metrics calculations, we have used 1× A100 40GB GPU (2× A100 for >30B models), cumulatively consuming approximately 3800 GPU-hours. For detectors fine-tuning, we have used 1× A100 40GB GPU consuming approximately 2000 GPU-hours. Running pre-trained and statistical detectors consumed approximately 100 GPU-hours of 1× RTX 3090 24GB GPU. For other tasks, we have not used GPU acceleration.
Appendix B Dataset Creation
Dataset preparation consisted of three important steps, namely selection of authentic human-written texts, machine-generation of texts, and final post-processing.
B.1 Human-Written Text Selection
Since no suitable multilingual and multi-platform social-media texts dataset was publicly available, we have combined human-written texts out of six existing multilingual datasets. Telegram data originated in Pushshift Telegram [^3] https://doi.org/10.5281/zenodo.3607497, containing 317M messages (Baumgartner et al., 2020). Twitter data originated in CLEF2022-CheckThat! Task 1444https://gitlab.com/checkthat_lab/clef2022-checkthat-lab/clef2022-checkthat-lab/-/tree/main/task1, containing 34k tweets on COVID-19 and politics (Nakov et al., 2022), combined with Sentiment140555https://www.kaggle.com/datasets/kazanova/sentiment140/data, containing 1.6M tweets on various topics (Go et al., 2009). Gab data originated in gab_posts_jan_2018666https://doi.org/10.5281/zenodo.1418347, containing 22M posts (Zannettou et al., 2018). Discord data originated in Discord-Data [^7] https://www.kaggle.com/datasets/jef1056/discord-data, containing 51M messages (Fan, 2021). And finally, WhatsApp data originated in whatsapp-public-groups [^8] https://github.com/gvrkiran/whatsapp-public-groups, containing 300k messages (Garimella and Tyson, 2018). These datasets have been deliberately selected due to containing older data (before 2022, most of them before 2020), when the text-generation AI have not been so mature in generation of multilingual texts, providing a higher confidence of the texts being actually written by humans (although cannot be 100% guaranteed).
The combined text samples have been deduplicated, resulting in over 283M texts, while using only texts with at least 3 words. We have used FastText [^9] https://pypi.org/project/fasttext language detection to get rough estimation for such a massive amount of texts (i.e., fast prediction with a reasonable accuracy), resulting in 176 different languages detected in the combined data. Based on such detected languages, we have pseudo-randomly sampled up to 10k texts for each available language from each of the five social-media platforms, resulting in about 2M of texts samples in the subset. Since social-media texts are quite short and often grammatically incorrect, the FastText language detection is quite noisy. Therefore, we have used four language detectors on the subset, namely FastText, Polyglot [^10] https://pypi.org/project/polyglot, Lingua [^11] https://pypi.org/project/lingua-language-detector, and LanguageIdentifier [^12] https://pypi.org/project/LanguageIdentifier.
To balance an accuracy of the language detection and minimization of unnecessary drop of samples, we have selected a combination of three-detectors match with a lower confidence predictions and two-detectors match with a higher confidence predictions, while removing URLs, hashtags, and user references in the texts for the detection purpose (the specific algorithm is provided in the source-code repository). Based on such a more accurate language detection, we have pseudo-randomly sampled up to 1300 texts (up to 300 for test split and the remaining up to 1000 for train split if available) for each of the selected 22 languages and platform. This process resulted in 61,592 human-written texts.
Approach METEOR $\uparrow$ BERTScore $\uparrow$ ngram $\uparrow$ LD $\downarrow$ MAUVE $\downarrow$ LangCheck $\downarrow$
k_to_one 0.163 (±0.33) 0.458 (±0.32) 0.108 (±0.18) 0.924 (±0.05) 0.148 35.18%
keywords 0.050 (±0.04) 0.537 (±0.16) 0.045 (±0.03) 1.973 (±0.81) 0.037 36.73%
paraphrase_1 0.439 (±0.13) 0.754 (±0.06) 0.322 (±0.15) 2.305 (±2.63) 0.336 24.32%
paraphrase_2 0.266 (±0.15) 0.682 (±0.07) 0.174 (±0.12) 4.123 (±6.11) 0.160 36.23%
paraphrase_3 0.209 (±0.12) 0.661 (±0.06) 0.133 (±0.10) 5.751 (±10.13) 0.130 37.59%
paraphrase_4 0.178 (±0.10) 0.647 (±0.06) 0.112 (±0.08) 7.627 (±14.68) 0.107 38.14%
paraphrase_5 0.151 (±0.09) 0.636 (±0.06) 0.092 (±0.07) 9.710 (±19.67) 0.095 38.81%
Table 9: Similarity analysis between machine-generated and human-written social-media texts subset of different approaches [mean ($\pm$ std)]. Arrows refer to values representing more similar texts, boldfaced values represent the most similar texts for each metric.
B.2 Social-Media Texts Generation
By using a small subset (10 samples per language) of the selected human-written texts, we have evaluated usability of multiple potential instruction-following LLMs for generation of social-media texts in the selected languages by using three different approaches, namely k-to-1 (10 human samples selected and used in a prompt for the model to generate a similar text, i.e., few-shot prompting), keywords (two longest words besides URLs and hashtags have been extracted from the human text and used in a prompt to be included in the generated text), and paraphrase (paraphrasing the text included in the prompt). Manual human check of the generated samples, revealed several problems of the approaches. The k-to-1 approach is sometimes not understandable for the models and we lose 1-to-1 mapping between human and machine samples. The keywords approach makes the generated text too different (out of context) from the original. On the other hand, the paraphrase approach makes the generated text too similar to the original. As shown in (Tripto et al., 2023), a single iteration of paraphrasing is not sufficient to confidently change the authorship (in our case from a human to a machine). Therefore, we have executed up to 5 iterations of paraphrasing and compared similarity metrics of different approaches (Table 9).
METEOR (Banerjee and Lavie, 2005) (used as a standard in machine translation) measures similarity based on unigrams. BERTScore (Zhang et al., 2019) with mBERT model measures contextual embeddings based similarity and is more robust to adversarial texts. ngram [^13] https://pypi.org/project/ngram (3-grams) is a language-independent string similarity metric in the form of a ratio of the shared ngrams between two strings. Higher values of these three metrics represent more similar texts. Levenshtein distance (LD) is used as a character-level edit distance [^14] https://github.com/roy-ht/editdistance, normalized to the text length, where a lower value represents more similar texts. MAUVE (Pillutla et al., 2021) score is used to measure a gap between distributions of human and machine texts. For the purpose of generation of similar texts, lower gap between distributions is better. LangCheck is a percentage of texts with changed languages based on FastText predictions.
However, we find MAUVE and LangCheck metrics as unreliable for such small amount of samples and lower text-lengths of social-media texts (providing them just for reference and comparison to metrics values of final full dataset). We have used longer and more formal (i.e., grammatically correct) news-domain texts to evaluate actual text-generation capability of the selected models in the selected languages (resulting in excluding Falcon-40B, Gemma-7B, Llama-2-13B from the selected generation models). Based on the automated similarity analysis and to balance cost-efficiency, we selected the 3 iteration of paraphrasing approach (also confirmed by Tripto et al., 2023 to converge towards paraphraser model authorship) for social-media text generation. The final prompt used for generation is as follows:
[…]
  Generate the text in {language_name}
  similar to the input social media text
  but using different words and sentence
[…]
  Output:
Generator METEOR $\uparrow$ BERTScore $\uparrow$ ngram $\uparrow$ LD $\downarrow$ MAUVE $\downarrow$ LangCheck $\downarrow$
aya-101 0.195 (±0.23) 0.675 (±0.14) 0.156 (±0.18) 1.104 (±0.91) 0.063 10.59%
gemini 0.152 (±0.20) 0.621 (±0.10) 0.087 (±0.17) 16.296 (±33.39) 0.025 6.16%
gpt-3.5-turbo-0125 0.143 (±0.18) 0.664 (±0.10) 0.080 (±0.10) 2.359 (±3.82) 0.076 22.80%
Mistral-7B-Instruct-v0.2 0.152 (±0.16) 0.652 (±0.08) 0.088 (±0.09) 2.383 (±1.93) 0.047 10.61%
opt-iml-max-30b 0.127 (±0.19) 0.659 (±0.12) 0.105 (±0.14) 0.998 (±0.57) 0.108 14.88%
v5-Eagle-7B-HF 0.108 (±0.15) 0.628 (±0.07) 0.071 (±0.10) 2.568 (±2.10) 0.027 6.01%
vicuna-13b 0.133 (±0.17) 0.650 (±0.09) 0.089 (±0.11) 1.811 (±1.40) 0.042 6.26%
Table 10: Similarity analysis between machine-generated (3-iteration paraphrased) and human-written (original) social-media texts for individual generation models [mean ($\pm$ std)]. Arrows refer to values representing more similar texts, boldfaced values represent the most similar texts for each metric.
Generator Empty Short Duplicate WC US UW
aya-101 1558 1821 2108 11.86 (±12.92) 0.97 (±0.16) 0.9 (±0.19)
gemini 16 126 36 71.08 (±54.01) 0.99 (±0.05) 0.73 (±0.16)
gpt-3.5-turbo-0125 13 34 2965 20.73 (±20.76) 1.0 (±0.02) 0.91 (±0.11)
Mistral-7B-Instruct-v0.2 14 110 85 18.76 (±14.21) 1.0 (±0.02) 0.89 (±0.11)
opt-iml-max-30b 1126 2197 1939 8.76 (±8.64) 0.98 (±0.14) 0.92 (±0.17)
v5-Eagle-7B-HF 15 287 28 22.13 (±17.17) 1.0 (±0.03) 0.88 (±0.11)
vicuna-13b 194 550 408 17.46 (±14.7) 1.0 (±0.06) 0.91 (±0.12)
human 0 3591 27 12.83 (±19.21) 1.0 (±0.01) 0.9 (±0.14)
Table 11: Statistics of the post-processed human-written and machine-generated social-media texts. WC refers to the word count, US refers to the unique sentences, and UW refers to the unique words [mean ($\pm$ std)].
We have set the $min\_new\_tokens$ to 5, $max\_new\_tokens$ to 200, $num\_return\_sequences$ to 1, using the nucleus sampling with $top\_p$ of 0.95 and $top\_k$ of 50 (not all of the parameters settable in API-based generation). After each paraphrasing iteration, the generated text is post-processed to remove redundant parts and to ensure the text is at most by 10 tokens longer than the original. We have used 3 trials to generate a paraphrase different than the original text, returning an empty string upon failure. Due to the safety filters in some of the LLMs (Gemini in the selected generation models), they tend to refuse generation of texts similar to “offensive” text present in some social-media texts. To limit generation failures, we have used a jailbreak151515a modified version of https://github.com/friuns2/BlackFriday-GPTs-Prompts/blob/main/gpts/evil-pure-evil.md for the research purpose.
Based on the similarity analysis of the final generated texts by the selected generation models (reported in Table 10), we can observe only small differences between the generators. Aya-101 generated the most similar texts in general, while OPT-IML-Max-30B provided the worst results (based on higher MAUVE and LangCheck scores). ChatGPT (GPT-3.5-Turbo) also generated the texts resulting in a little higher MAUVE score and the highest language mismatch (LangCheck). However, the language mismatch of ChatGPT (using all the selected languages) was not confirmed by longer news articles generation (FastText language detection in such texts is definitely more accurate), where it actually achieved the lowest (under 1%) LangCheck score among the generators. We assume that in social-media text generation it can better follow the original style of the text than the other generators (e.g., grammatically incorrect, slang), which is more difficult for accurate language detection.
B.3 Post-processing of Generated Texts
Both, the human and machine texts are cleaned by removing leading and trailing white-spaces, removing characters making problems in Polyglot161616based on https://github.com/aboSamoor/polyglot/issues/71#issuecomment-707997790, truncating the parts of texts above 200 words, and dropping duplicates and the samples with less than 3 words. Thus, we ensure that no text sample has multiple labels and that both human and machine samples are processed in the same way.
The linguistic-analysis statistics of the post-processed texts are provided in Table 11. Based on the statistics, we can see that Aya-101 and OPT-IML-Max-30B failed to generate a paraphrase in higher amounts of texts, also generating shorter texts than the others. On the other hand, Gemini generated the longest texts, but also has the lowest ratio of unique words (inferring higher repetitiveness, maybe connected to the integrated safety filters in spite of using a jailbreak). Also, a higher amount of human texts have not contained enough (at least 3) words after post-processing. ChatGPT (GPT-3.5-Turbo) generated the highest amount of duplicates. These filtered amounts however not affected well-balancing of the dataset across generators (ranging from 56k samples for OPT-IML to 61.3k samples for Mistral), resulting in the final MultiSocial dataset of 472,097 texts (of which about 58k are human-written). Regarding the final composition of the dataset across languages and platforms, we provide the sample counts in Table 2.
We have even conducted manual human check of the generated texts, resulting in identification of various phrases indicating noise in the data that has not been removed by the post-processing, such as “as an ai model”, “language model”, “instruction”, “task”, etc. After a deeper analysis, such texts are present in about 1% of the data. We are leaving these texts in the MultiSocial dataset for further analysis purposes (e.g., analysis of model failures across generators and across languages); however, they are filtered-out in the pre-processing step of our experiments. The identified noisy text samples are clearly marked in the published dataset.
Appendix C Fine-tuning Settings
For the fine-tuning process of the fine-tuned detection methods, we have used a parameter efficient fine-tuning (PEFT) technique called QLoRA (Dettmers et al., 2023) with default parameters (except for target_modules set to query_key_value and $r=4$). The training process used the AdamW optimizer with the linear scheduler and the learning rate of $2E^{-4}$. Batch size of 2 with gradient accumulation steps of 8 have been used. We have used fixed 1 epoch for training (7 epochs in case of smaller subset selection), but also limiting the models training process to 48 hours (checkpointing each 20% of the epoch). All the settings can be found in the source code available in the published repository, enabling full replication. We aimed to use the same training settings across the various detection models training; however, we have used full fine-tuning for the XLM-RoBERTa-large model instead of QLoRA due to lack of support.
Due to a high class imbalance when using multiple generators data, we have experimented with various class balancing strategies for training. Namely, no balancing, majority-class downsampling, minority-class upsampling, mixed up-down-sampling (duplicating the minority-class samples just once and afterwards downsampling the majority-class). The results (using accuracy and AUC ROC metrics) indicated that there is a negligible effect on performance based on the used strategy, while no balancing having slower learning ability (however, the performance is eventually competitive with the others). Since having no significant impact on the performance, we have used majority-class downsampling to limit the number of steps in epoch.
Appendix D Cross-domain Evaluation
For the evaluation on the out-of-domain data, we use the news articles of MULTITuDE (Macko et al., 2023) benchmark. We have used the published scripts [^17] https://github.com/kinit-sk/mgt-detection-benchmark to extend the test set to our selection of languages. For the text generation, we have used the same models as in the proposed MultiSocial dataset (to evaluate only cross-domain capability), while we used Llama-2-70b model instead for Gemini for out-of-distribution (cross-generator) evaluation.
In Table 12 and 13, the results are provided in the same way as in Table 3, while testing on MULTITuDE (news domain) data. In pure cross-domain evaluation (Table 12), the machine texts are generated by the same generators (as used for training), in out-of-distribution evaluation (Table 13), the machine texts are generated only by Llama-2-70b (not available in MultiSocial for training).
Rank Detector AUC ROC
MacroF1
@5%FPR
1 Llama-3-8b-MultiSocial 0.9273 0.7988
2 Aya-101-MultiSocial 0.9262 0.8008
3 mDeBERTa-v3-base-MultiSocial 0.9025 0.7512
4 Mistral-7b-v0.1-MultiSocial 0.8988 0.7937
5 XLM-RoBERTa-large-MultiSocial 0.8309 0.7306
6 Binoculars 0.8303 0.4041
7 Fast-Detect-GPT 0.8104 0.6361
8 Falcon-rw-1b-MultiSocial 0.7592 0.6394
9 BLOOMZ-3b-MultiSocial 0.7071 0.5731
10 LLM-Deviation 0.6568 0.3568
11 DetectLLM-LRR 0.6496 0.4133
12 S5 0.6336 0.3519
13 Longformer Detector 0.6157 0.2564
14 ChatGPT-Detector-RoBERTa-Chinese 0.5896 0.4296
15 RoBERTa-large-OpenAI-Detector 0.5707 0.1958
16 BLOOMZ-3b-mixed-Detector 0.5536 0.1891
17 ruRoBERTa-ruatd-binary 0.5186 0.1485
Table 12: Cross-domain evaluation of the selected MGTD methods of statistical, pre-trained, and fine-tuned categories.
Rank Detector AUC ROC
MacroF1
@5%FPR
1 Fast-Detect-GPT 0.9238 0.8471
2 Binoculars 0.9048 0.7568
3 mDeBERTa-v3-base-MultiSocial 0.8871 0.8011
4 Mistral-7b-v0.1-MultiSocial 0.8614 0.7673
5 Aya-101-MultiSocial 0.8574 0.7556
6 Llama-3-8b-MultiSocial 0.8549 0.7352
7 XLM-RoBERTa-large-MultiSocial 0.7928 0.7108
8 Falcon-rw-1b-MultiSocial 0.7710 0.6912
9 DetectLLM-LRR 0.7559 0.7121
10 LLM-Deviation 0.7257 0.5931
11 Longformer Detector 0.7032 0.5018
12 S5 0.6849 0.5506
13 ChatGPT-Detector-RoBERTa-Chinese 0.6788 0.6161
14 BLOOMZ-3b-MultiSocial 0.6515 0.6026
15 RoBERTa-large-OpenAI-Detector 0.5596 0.4315
16 ruRoBERTa-ruatd-binary 0.5327 0.3473
17 BLOOMZ-3b-mixed-Detector 0.5212 0.4096
Table 13: Out-of-distribution (cross-domain and cross-generator) evaluation of the selected MGTD methods of statistical, pre-trained, and fine-tuned categories.
Appendix E Ablation Study
Train Test Language [AUC ROC mean]
Language ar bg ca cs de el en es et ga gd hr hu nl pl pt ro ru sk sl uk zh all
en 0.78 0.86 0.77 0.87 0.88 0.88 0.96 0.85 0.89 N/A N/A 0.92 0.95 0.82 0.86 0.90 0.92 0.85 N/A N/A 0.79 0.65 0.84
es 0.78 0.84 0.83 0.83 0.88 0.84 0.87 0.94 0.86 N/A N/A 0.88 0.92 0.81 0.84 0.91 0.91 0.86 N/A N/A 0.79 0.64 0.82
ru 0.73 0.90 0.68 0.81 0.81 0.84 0.83 0.76 0.82 N/A N/A 0.84 0.86 0.74 0.82 0.81 0.83 0.93 N/A N/A 0.87 0.61 0.77
en-es-ru 0.88 0.92 0.86 0.91 0.91 0.93 0.95 0.94 0.91 N/A N/A 0.93 0.96 0.84 0.89 0.93 0.93 0.93 N/A N/A 0.88 0.84 0.90
Table 14: Cross-lingual mean AUC ROC performance of the MGT detectors with autoregressive foundational models fine-tuned monolingually (en, es and ru) and multilingually (en-es-ru), evaluated based on Telegram data (for training as well as for testing). N/A refers to not enough samples (at least 2000) in MultiSocial Telegram data.
Train Test Language [AUC ROC mean]
Language ar bg ca cs de el en es et ga gd hr hu nl pl pt ro ru sk sl uk zh all
en 0.85 0.95 0.79 0.96 0.88 0.93 0.96 0.90 0.95 N/A N/A 0.96 0.99 0.86 0.93 0.94 0.96 0.90 N/A N/A 0.85 0.83 0.91
es 0.88 0.96 0.86 0.96 0.91 0.92 0.94 0.94 0.95 N/A N/A 0.96 0.99 0.87 0.93 0.94 0.96 0.92 N/A N/A 0.87 0.86 0.92
ru 0.91 0.97 0.86 0.96 0.88 0.94 0.94 0.90 0.96 N/A N/A 0.96 0.98 0.86 0.94 0.93 0.95 0.96 N/A N/A 0.92 0.90 0.93
en-es-ru 0.91 0.97 0.87 0.96 0.91 0.93 0.96 0.93 0.96 N/A N/A 0.96 0.99 0.88 0.94 0.95 0.97 0.94 N/A N/A 0.90 0.89 0.93
Table 15: Cross-lingual mean AUC ROC performance of the MGT detectors with non-autoregressive foundational models fine-tuned monolingually (en, es and ru) and multilingually (en-es-ru), evaluated based on Telegram data (for training as well as for testing). N/A refers to not enough samples (at least 2000) in MultiSocial Telegram data.
Based on Table 25, we have identified two groups of detectors based on their performances across languages, namely autoregressive models and others. Autoregressive group includes Llama-3-8b, Mistral-7b-v0.1, BLOOMZ-3b, and Falcon-rw-1b. The non-autoregressive group includes Aya-101, XLM-RoBERTa-large, and mDeBERTa-v3-base. The summarized results, analogous to the Table 7, but provided for each group separately in Table 14 and Table 15. There are clear differences between the results of these groups, since in case of non-autogregressive group, we can see no significant differences between monolingually and multilingually fine-tuned detectors.
Another possible explanation of such a different behavior of some detectors is their pre-training on a huge number of languages (100+ languages in case of each detector in the non-autoregressive group). However, the BLOOMZ model has also been pre-trained on 50+ languages, and still shows a behavior similar to others in the autoregressive group.
Appendix F Results Data
Tables 16-18 contain per-generator AUC ROC performance of each MGT detection method for each test language, Tables 19-21 contain per-platform AUC ROC performance of each MGT detection method for each test language, and Tables 22-24 contains per-generator AUC ROC performance of each MGT detection method for each platform, separately for each MGTD category. Table 25 contains cross-lingual evaluation of differently fine-tuned MGT detectors. Table 26 contains cross-platform evaluation of differently fine-tuned MGT detectors.
Test Language [AUC ROC]
Detector Generator ar bg ca cs de el en es et ga gd hr hu nl pl pt ro ru sk sl uk zh all
Binoculars Aya-101 0.70 0.61 0.60 0.69 0.75 0.77 0.79 0.73 0.67 0.62 0.59 0.72 0.71 0.71 0.71 0.70 0.67 0.67 0.64 0.67 0.54 0.65 0.69
GPT-3.5-Turbo-0125 0.62 0.64 0.57 0.72 0.72 0.78 0.75 0.70 0.66 0.61 0.59 0.75 0.74 0.68 0.73 0.70 0.71 0.69 0.70 0.66 0.61 0.92 0.68
Gemini 0.64 0.73 0.71 0.85 0.86 0.80 0.92 0.90 0.87 0.95 N/A 0.88 0.87 0.80 0.84 0.88 0.88 0.69 0.82 0.85 0.70 0.90 0.83
Mistral-7B-Instruct-v0.2 0.69 0.67 0.61 0.70 0.67 0.75 0.71 0.69 0.72 0.73 0.82 0.78 0.75 0.64 0.71 0.70 0.72 0.70 0.66 0.68 0.65 0.63 0.68
OPT-IML-Max-30b 0.77 0.65 0.58 0.58 0.64 0.75 0.74 0.64 0.68 0.64 0.63 0.71 0.73 0.63 0.62 0.66 0.63 0.65 0.58 0.65 0.64 0.43 0.64
Vicuna-13b 0.77 0.69 0.63 0.79 0.80 0.86 0.82 0.78 0.75 0.76 0.84 0.82 0.80 0.78 0.81 0.80 0.79 0.80 0.76 0.76 0.64 0.76 0.76
v5-Eagle-7B-HF 0.73 0.72 0.64 0.82 0.82 0.84 0.89 0.85 0.79 0.82 0.80 0.83 0.84 0.81 0.81 0.82 0.81 0.81 0.80 0.81 0.70 0.86 0.79
DetectLLM-LRR Aya-101 0.75 0.82 0.67 0.89 0.75 0.84 0.77 0.76 0.82 0.73 0.69 0.79 0.91 0.76 0.82 0.76 0.81 0.74 0.79 0.70 0.66 0.66 0.70
GPT-3.5-Turbo-0125 0.70 0.75 0.65 0.93 0.74 0.83 0.78 0.77 0.85 0.78 0.65 0.88 0.92 0.77 0.85 0.80 0.84 0.68 0.86 0.77 0.66 0.76 0.71
Gemini 0.89 0.94 0.65 0.96 0.86 0.95 0.90 0.94 0.96 0.86 N/A 0.97 0.97 0.84 0.92 0.93 0.96 0.84 0.94 0.93 0.86 0.92 0.83
Mistral-7B-Instruct-v0.2 0.75 0.85 0.68 0.92 0.72 0.85 0.76 0.80 0.86 0.82 0.87 0.88 0.93 0.70 0.85 0.84 0.84 0.77 0.79 0.74 0.75 0.72 0.71
OPT-IML-Max-30b 0.68 0.81 0.63 0.84 0.66 0.83 0.70 0.70 0.83 0.63 0.51 0.80 0.89 0.66 0.82 0.73 0.79 0.66 0.75 0.64 0.62 0.60 0.67
Vicuna-13b 0.90 0.90 0.72 0.96 0.85 0.92 0.81 0.87 0.89 0.88 0.83 0.91 0.95 0.86 0.93 0.90 0.91 0.90 0.88 0.82 0.79 0.84 0.78
v5-Eagle-7B-HF 0.87 0.92 0.79 0.97 0.87 0.93 0.88 0.92 0.95 0.86 0.86 0.94 0.97 0.90 0.95 0.92 0.95 0.88 0.94 0.91 0.86 0.92 0.82
Fast-Detect-GPT Aya-101 0.75 0.66 0.64 0.83 0.81 0.74 0.79 0.75 0.75 0.73 0.71 0.74 0.79 0.78 0.78 0.76 0.77 0.72 0.75 0.75 0.67 0.68 0.75
GPT-3.5-Turbo-0125 0.74 0.61 0.54 0.82 0.71 0.72 0.73 0.64 0.69 0.71 0.73 0.79 0.70 0.70 0.74 0.71 0.70 0.73 0.75 0.75 0.73 0.80 0.71
Gemini 0.86 0.85 0.70 0.88 0.88 0.89 0.92 0.88 0.90 0.88 N/A 0.92 0.89 0.81 0.84 0.89 0.88 0.82 0.88 0.90 0.83 0.90 0.87
Mistral-7B-Instruct-v0.2 0.60 0.50 0.51 0.66 0.56 0.43 0.69 0.58 0.47 0.58 0.80 0.66 0.63 0.54 0.60 0.61 0.57 0.59 0.46 0.55 0.56 0.58 0.58
OPT-IML-Max-30b 0.64 0.50 0.58 0.72 0.70 0.48 0.77 0.69 0.65 0.62 0.54 0.76 0.70 0.67 0.74 0.73 0.74 0.57 0.65 0.59 0.50 0.52 0.67
Vicuna-13b 0.81 0.67 0.65 0.83 0.85 0.57 0.83 0.80 0.52 0.75 0.79 0.83 0.79 0.82 0.84 0.84 0.83 0.83 0.57 0.77 0.74 0.78 0.78
v5-Eagle-7B-HF 0.85 0.79 0.72 0.91 0.86 0.80 0.89 0.85 0.85 0.78 0.84 0.88 0.86 0.85 0.87 0.86 0.89 0.85 0.88 0.85 0.81 0.91 0.85
LLM-Deviation Aya-101 0.79 0.84 0.66 0.90 0.77 0.84 0.76 0.76 0.85 0.76 0.73 0.81 0.91 0.77 0.83 0.77 0.82 0.75 0.81 0.73 0.67 0.69 0.70
GPT-3.5-Turbo-0125 0.68 0.73 0.63 0.93 0.74 0.83 0.77 0.76 0.86 0.79 0.68 0.88 0.92 0.76 0.86 0.80 0.84 0.66 0.87 0.79 0.65 0.78 0.71
Gemini 0.90 0.95 0.64 0.96 0.87 0.96 0.90 0.94 0.97 0.86 N/A 0.97 0.98 0.85 0.93 0.93 0.96 0.85 0.94 0.95 0.88 0.92 0.82
Mistral-7B-Instruct-v0.2 0.76 0.82 0.66 0.92 0.72 0.86 0.75 0.79 0.86 0.85 0.93 0.89 0.93 0.68 0.85 0.83 0.84 0.74 0.78 0.74 0.72 0.70 0.70
OPT-IML-Max-30b 0.77 0.85 0.62 0.86 0.67 0.84 0.70 0.71 0.84 0.69 0.58 0.83 0.91 0.67 0.83 0.75 0.81 0.71 0.77 0.68 0.66 0.67 0.68
Vicuna-13b 0.92 0.91 0.72 0.96 0.86 0.93 0.81 0.86 0.90 0.89 0.88 0.91 0.96 0.86 0.94 0.90 0.92 0.89 0.88 0.83 0.79 0.84 0.78
v5-Eagle-7B-HF 0.89 0.93 0.79 0.98 0.87 0.94 0.88 0.93 0.97 0.88 0.89 0.95 0.98 0.91 0.96 0.93 0.95 0.88 0.95 0.92 0.87 0.92 0.82
S5 Aya-101 0.80 0.84 0.67 0.89 0.76 0.85 0.75 0.75 0.84 0.75 0.71 0.81 0.90 0.76 0.82 0.77 0.82 0.75 0.80 0.73 0.67 0.70 0.71
GPT-3.5-Turbo-0125 0.67 0.71 0.62 0.92 0.73 0.82 0.74 0.73 0.84 0.77 0.67 0.87 0.91 0.74 0.84 0.78 0.84 0.65 0.85 0.77 0.65 0.74 0.70
Gemini 0.84 0.94 0.63 0.95 0.85 0.95 0.88 0.92 0.96 0.84 N/A 0.97 0.97 0.83 0.91 0.92 0.96 0.83 0.93 0.93 0.87 0.90 0.81
Mistral-7B-Instruct-v0.2 0.72 0.79 0.66 0.91 0.71 0.85 0.72 0.78 0.83 0.83 0.92 0.88 0.92 0.67 0.83 0.81 0.84 0.72 0.76 0.72 0.70 0.66 0.69
OPT-IML-Max-30b 0.79 0.84 0.63 0.86 0.67 0.84 0.69 0.71 0.84 0.71 0.63 0.82 0.90 0.67 0.83 0.74 0.81 0.73 0.77 0.69 0.66 0.71 0.69
Vicuna-13b 0.91 0.89 0.72 0.95 0.85 0.93 0.79 0.85 0.89 0.88 0.86 0.90 0.95 0.85 0.93 0.90 0.91 0.88 0.87 0.82 0.79 0.82 0.78
v5-Eagle-7B-HF 0.88 0.92 0.79 0.97 0.86 0.94 0.86 0.91 0.96 0.86 0.85 0.94 0.97 0.89 0.95 0.92 0.95 0.88 0.95 0.90 0.86 0.91 0.81
Table 16: Per-LLM AUC ROC performance of statistical MGT detectors category. N/A refers to not enough samples per each class (at least 10) making AUC ROC value irrelevant.
Test Language [AUC ROC]
Detector Generator ar bg ca cs de el en es et ga gd hr hu nl pl pt ro ru sk sl uk zh all
BLOOMZ-3b-mixed-Detector Aya-101 0.83 0.88 0.82 0.88 0.85 0.82 0.85 0.84 0.91 0.79 0.68 0.86 0.92 0.84 0.84 0.85 0.78 0.78 0.83 0.76 0.81 0.86 0.83
GPT-3.5-Turbo-0125 0.85 0.80 0.81 0.86 0.82 0.82 0.87 0.83 0.87 0.81 0.67 0.83 0.87 0.80 0.83 0.83 0.74 0.73 0.80 0.71 0.82 0.19 0.80
Gemini 0.64 0.55 0.63 0.69 0.65 0.76 0.50 0.58 0.78 0.65 N/A 0.67 0.77 0.66 0.72 0.63 0.33 0.48 0.65 0.41 0.43 0.60 0.59
Mistral-7B-Instruct-v0.2 0.76 0.71 0.74 0.74 0.71 0.71 0.85 0.78 0.80 0.75 0.57 0.75 0.78 0.68 0.73 0.78 0.72 0.68 0.68 0.59 0.71 0.76 0.74
OPT-IML-Max-30b 0.69 0.66 0.76 0.80 0.71 0.72 0.78 0.79 0.77 0.71 0.60 0.76 0.77 0.71 0.75 0.77 0.70 0.63 0.79 0.69 0.65 0.75 0.73
Vicuna-13b 0.88 0.77 0.84 0.79 0.81 0.74 0.88 0.83 0.81 0.78 0.77 0.77 0.85 0.78 0.81 0.83 0.66 0.75 0.78 0.67 0.73 0.83 0.79
v5-Eagle-7B-HF 0.86 0.82 0.85 0.82 0.82 0.76 0.90 0.86 0.89 0.83 0.65 0.79 0.88 0.79 0.81 0.86 0.72 0.72 0.77 0.67 0.74 0.85 0.81
ChatGPT-Detector-RoBERTa-Chinese Aya-101 0.75 0.79 0.73 0.63 0.75 0.70 0.82 0.75 0.78 0.61 0.52 0.61 0.72 0.67 0.61 0.61 0.67 0.73 0.60 0.60 0.71 0.76 0.67
GPT-3.5-Turbo-0125 0.59 0.63 0.64 0.59 0.70 0.65 0.86 0.73 0.76 0.65 0.52 0.57 0.68 0.63 0.57 0.69 0.68 0.62 0.56 0.60 0.63 0.88 0.66
Gemini 0.74 0.84 0.80 0.78 0.88 0.72 0.97 0.90 0.94 0.95 N/A 0.76 0.89 0.78 0.71 0.86 0.90 0.81 0.73 0.78 0.87 0.78 0.80
Mistral-7B-Instruct-v0.2 0.73 0.77 0.79 0.69 0.82 0.78 0.92 0.84 0.84 0.87 0.66 0.73 0.82 0.73 0.63 0.84 0.84 0.76 0.66 0.67 0.76 0.75 0.76
OPT-IML-Max-30b 0.79 0.84 0.74 0.60 0.70 0.76 0.82 0.74 0.73 0.65 0.58 0.52 0.68 0.63 0.60 0.62 0.59 0.74 0.62 0.60 0.70 0.63 0.65
Vicuna-13b 0.74 0.86 0.81 0.63 0.82 0.80 0.93 0.82 0.74 0.73 0.73 0.60 0.81 0.70 0.62 0.73 0.76 0.82 0.62 0.68 0.72 0.88 0.72
v5-Eagle-7B-HF 0.74 0.87 0.81 0.71 0.88 0.81 0.95 0.89 0.88 0.83 0.76 0.59 0.83 0.72 0.63 0.77 0.78 0.83 0.66 0.65 0.80 0.95 0.76
Longformer Detector Aya-101 0.22 0.37 0.25 0.36 0.32 0.41 0.53 0.34 0.35 0.38 0.36 0.39 0.35 0.29 0.35 0.31 0.31 0.36 0.35 0.37 0.43 0.27 0.36
GPT-3.5-Turbo-0125 0.39 0.50 0.37 0.42 0.44 0.50 0.59 0.48 0.40 0.39 0.43 0.48 0.46 0.43 0.44 0.40 0.42 0.45 0.43 0.43 0.55 0.54 0.45
Gemini 0.67 0.61 0.47 0.42 0.46 0.82 0.80 0.48 0.50 0.92 N/A 0.52 0.49 0.44 0.46 0.42 0.36 0.64 0.34 0.33 0.80 0.76 0.53
Mistral-7B-Instruct-v0.2 0.33 0.48 0.37 0.58 0.51 0.54 0.63 0.50 0.61 0.64 0.53 0.61 0.62 0.49 0.56 0.51 0.49 0.47 0.59 0.55 0.66 0.50 0.52
OPT-IML-Max-30b 0.17 0.42 0.28 0.33 0.32 0.39 0.50 0.33 0.41 0.44 0.38 0.42 0.43 0.31 0.35 0.35 0.35 0.29 0.33 0.44 0.51 0.26 0.36
Vicuna-13b 0.28 0.47 0.24 0.55 0.42 0.53 0.66 0.42 0.48 0.58 0.50 0.54 0.50 0.40 0.49 0.41 0.45 0.49 0.51 0.50 0.60 0.42 0.46
v5-Eagle-7B-HF 0.31 0.49 0.30 0.61 0.49 0.58 0.81 0.50 0.60 0.67 0.68 0.63 0.61 0.47 0.56 0.51 0.48 0.52 0.59 0.56 0.68 0.54 0.53
RoBERTa-large-OpenAI-Detector Aya-101 0.76 0.46 0.46 0.19 0.40 0.72 0.73 0.43 0.32 0.37 0.46 0.31 0.20 0.39 0.31 0.35 0.36 0.57 0.35 0.45 0.43 0.69 0.43
GPT-3.5-Turbo-0125 0.60 0.41 0.45 0.16 0.34 0.59 0.48 0.34 0.25 0.28 0.51 0.20 0.17 0.31 0.23 0.26 0.28 0.48 0.27 0.38 0.30 0.61 0.35
Gemini 0.65 0.30 0.51 0.05 0.14 0.76 0.14 0.11 0.05 0.41 N/A 0.06 0.04 0.13 0.10 0.08 0.05 0.47 0.08 0.08 0.29 0.43 0.21
Mistral-7B-Instruct-v0.2 0.69 0.39 0.38 0.10 0.30 0.72 0.40 0.25 0.22 0.14 0.30 0.16 0.11 0.30 0.19 0.17 0.21 0.49 0.26 0.34 0.30 0.55 0.31
OPT-IML-Max-30b 0.86 0.60 0.54 0.30 0.52 0.79 0.71 0.49 0.32 0.51 0.67 0.34 0.24 0.51 0.36 0.40 0.40 0.70 0.48 0.54 0.60 0.83 0.49
Vicuna-13b 0.80 0.45 0.40 0.10 0.29 0.79 0.57 0.30 0.17 0.21 0.47 0.18 0.12 0.26 0.17 0.20 0.23 0.54 0.21 0.31 0.36 0.60 0.33
v5-Eagle-7B-HF 0.78 0.40 0.34 0.08 0.28 0.83 0.62 0.24 0.11 0.33 0.48 0.13 0.07 0.20 0.15 0.16 0.21 0.52 0.18 0.26 0.26 0.52 0.31
ruRoBERTa-ruatd-binary Aya-101 0.50 0.61 0.57 0.47 0.48 0.45 0.60 0.54 0.48 0.47 0.49 0.53 0.56 0.57 0.48 0.50 0.43 0.64 0.52 0.57 0.50 0.45 0.53
GPT-3.5-Turbo-0125 0.46 0.66 0.57 0.45 0.46 0.39 0.58 0.50 0.45 0.46 0.45 0.51 0.50 0.52 0.45 0.45 0.41 0.69 0.48 0.48 0.52 0.22 0.50
Gemini 0.17 0.25 0.53 0.14 0.19 0.11 0.19 0.17 0.11 0.30 N/A 0.10 0.12 0.24 0.17 0.15 0.08 0.30 0.12 0.11 0.23 0.26 0.18
Mistral-7B-Instruct-v0.2 0.42 0.69 0.53 0.50 0.49 0.35 0.58 0.52 0.49 0.45 0.38 0.57 0.53 0.56 0.48 0.49 0.41 0.79 0.52 0.52 0.69 0.52 0.53
OPT-IML-Max-30b 0.64 0.82 0.63 0.54 0.52 0.50 0.66 0.55 0.54 0.50 0.44 0.57 0.60 0.59 0.52 0.53 0.43 0.92 0.61 0.65 0.91 0.58 0.59
Vicuna-13b 0.31 0.69 0.54 0.45 0.43 0.36 0.66 0.49 0.43 0.51 0.57 0.53 0.49 0.51 0.44 0.45 0.34 0.82 0.45 0.50 0.62 0.57 0.51
v5-Eagle-7B-HF 0.35 0.72 0.53 0.50 0.44 0.29 0.66 0.51 0.51 0.59 0.50 0.56 0.50 0.51 0.51 0.46 0.33 0.78 0.45 0.50 0.66 0.45 0.52
Table 17: Per-LLM AUC ROC performance of pre-trained MGT detectors category. N/A refers to not enough samples per each class (at least 10) making AUC ROC value irrelevant.
Test Language [AUC ROC]
Detector Generator ar bg ca cs de el en es et ga gd hr hu nl pl pt ro ru sk sl uk zh all
Aya-101-MultiSocial Aya-101 0.92 0.98 0.97 0.98 0.97 0.96 0.96 0.96 0.98 0.92 0.91 0.98 0.99 0.97 0.97 0.96 0.96 0.93 0.97 0.92 0.90 0.93 0.96
GPT-3.5-Turbo-0125 0.99 1.00 1.00 0.99 0.99 0.99 1.00 0.99 0.99 0.97 0.96 0.99 1.00 0.99 0.99 0.99 0.99 0.99 0.99 0.97 0.98 0.99 0.99
Gemini 0.90 0.96 0.88 0.95 0.92 0.92 0.96 0.96 0.99 0.85 N/A 0.96 0.99 0.92 0.94 0.94 0.96 0.83 0.92 0.89 0.87 0.94 0.93
Mistral-7B-Instruct-v0.2 0.99 1.00 0.99 1.00 0.99 0.99 1.00 0.99 0.99 0.99 0.99 0.99 1.00 0.98 0.99 0.99 1.00 0.99 1.00 0.98 0.98 0.98 0.99
OPT-IML-Max-30b 0.98 0.99 0.96 0.97 0.94 0.96 0.92 0.95 0.95 0.84 0.77 0.95 0.98 0.95 0.97 0.95 0.96 0.98 0.96 0.91 0.95 0.99 0.95
Vicuna-13b 1.00 1.00 0.99 1.00 0.99 0.99 1.00 0.99 0.99 0.98 0.95 0.99 1.00 0.99 0.99 0.99 1.00 0.99 0.99 0.98 0.97 0.99 0.99
v5-Eagle-7B-HF 0.99 1.00 1.00 1.00 0.99 0.99 1.00 1.00 1.00 0.98 0.94 1.00 1.00 0.99 1.00 1.00 1.00 0.99 1.00 0.98 0.99 0.99 0.99
BLOOMZ-3b-MultiSocial Aya-101 0.92 0.97 0.96 0.96 0.95 0.94 0.96 0.96 0.96 0.86 0.71 0.94 0.99 0.94 0.94 0.95 0.92 0.90 0.94 0.84 0.85 0.90 0.94
GPT-3.5-Turbo-0125 0.99 0.99 0.99 0.98 0.98 0.97 0.99 0.99 0.98 0.90 0.84 0.97 0.99 0.97 0.97 0.99 0.96 0.97 0.97 0.91 0.93 0.99 0.98
Gemini 0.91 0.93 0.88 0.96 0.91 0.93 0.97 0.97 0.99 0.91 N/A 0.95 0.99 0.90 0.92 0.96 0.93 0.82 0.89 0.75 0.77 0.96 0.93
Mistral-7B-Instruct-v0.2 0.96 0.99 0.98 0.98 0.97 0.97 1.00 0.98 0.99 0.95 0.81 0.98 1.00 0.93 0.96 0.99 0.97 0.96 0.98 0.91 0.94 0.98 0.97
OPT-IML-Max-30b 0.95 0.98 0.93 0.94 0.89 0.93 0.93 0.94 0.94 0.84 0.76 0.90 0.97 0.89 0.92 0.93 0.90 0.96 0.94 0.86 0.92 0.97 0.93
Vicuna-13b 0.99 0.99 0.98 0.98 0.98 0.98 1.00 0.99 0.98 0.93 0.88 0.97 0.99 0.97 0.98 0.99 0.96 0.97 0.98 0.90 0.91 0.99 0.98
v5-Eagle-7B-HF 0.98 0.99 0.99 0.99 0.98 0.98 1.00 1.00 0.99 0.94 0.86 0.99 1.00 0.98 0.98 1.00 0.99 0.98 0.98 0.95 0.95 1.00 0.99
Falcon-rw-1b-MultiSocial Aya-101 0.93 0.97 0.97 0.97 0.95 0.94 0.96 0.96 0.97 0.88 0.82 0.95 0.99 0.95 0.95 0.95 0.93 0.91 0.95 0.85 0.87 0.91 0.94
GPT-3.5-Turbo-0125 0.98 0.99 0.99 0.98 0.98 0.98 0.99 0.98 0.98 0.93 0.89 0.98 0.99 0.98 0.98 0.98 0.97 0.97 0.97 0.91 0.94 0.99 0.98
Gemini 0.86 0.96 0.89 0.95 0.92 0.93 0.99 0.94 0.98 0.92 N/A 0.95 0.99 0.91 0.92 0.92 0.92 0.87 0.88 0.72 0.87 0.90 0.93
Mistral-7B-Instruct-v0.2 0.95 0.99 0.98 0.98 0.97 0.97 1.00 0.97 0.99 0.96 0.93 0.98 0.99 0.95 0.97 0.98 0.98 0.94 0.97 0.89 0.92 0.95 0.97
OPT-IML-Max-30b 0.96 0.98 0.94 0.95 0.90 0.95 0.93 0.93 0.94 0.84 0.72 0.92 0.98 0.90 0.93 0.92 0.92 0.95 0.95 0.88 0.91 0.97 0.93
Vicuna-13b 0.98 0.98 0.98 0.99 0.98 0.98 1.00 0.98 0.99 0.95 0.93 0.97 0.99 0.97 0.98 0.98 0.97 0.96 0.98 0.89 0.88 0.98 0.98
v5-Eagle-7B-HF 0.98 0.99 0.98 0.99 0.98 0.98 1.00 0.99 0.99 0.96 0.92 0.99 1.00 0.98 0.98 0.99 0.99 0.97 0.98 0.94 0.95 0.99 0.98
Llama-3-8b-MultiSocial Aya-101 0.94 0.98 0.98 0.99 0.97 0.96 0.97 0.97 0.98 0.90 0.81 0.98 0.99 0.97 0.97 0.96 0.95 0.94 0.97 0.91 0.92 0.95 0.96
GPT-3.5-Turbo-0125 0.99 1.00 1.00 1.00 0.99 0.99 1.00 0.99 0.99 0.95 0.93 0.99 1.00 0.99 0.99 0.99 0.99 0.99 0.99 0.96 0.98 1.00 0.99
Gemini 0.89 0.96 0.91 0.96 0.95 0.89 0.99 0.97 0.98 0.91 N/A 0.98 0.99 0.94 0.96 0.96 0.96 0.87 0.94 0.91 0.88 0.96 0.95
Mistral-7B-Instruct-v0.2 0.99 1.00 0.99 0.99 0.99 0.99 1.00 0.99 0.99 0.99 0.99 0.99 1.00 0.98 0.99 0.99 0.99 0.99 0.99 0.97 0.98 0.99 0.99
OPT-IML-Max-30b 0.98 0.99 0.97 0.98 0.94 0.96 0.94 0.96 0.96 0.85 0.72 0.96 0.99 0.95 0.97 0.96 0.95 0.98 0.97 0.92 0.96 0.99 0.96
Vicuna-13b 1.00 1.00 0.99 1.00 0.99 0.99 1.00 0.99 0.99 0.98 0.92 0.99 1.00 0.99 0.99 0.99 0.99 0.99 0.99 0.96 0.97 1.00 0.99
v5-Eagle-7B-HF 0.99 1.00 1.00 1.00 0.99 0.99 1.00 1.00 1.00 0.98 0.98 1.00 1.00 0.99 1.00 1.00 1.00 0.99 0.99 0.98 0.99 1.00 0.99
Mistral-7b-v0.1-MultiSocial Aya-101 0.94 0.98 0.97 0.98 0.97 0.95 0.97 0.97 0.98 0.90 0.89 0.98 0.99 0.97 0.97 0.96 0.95 0.94 0.96 0.90 0.91 0.94 0.96
GPT-3.5-Turbo-0125 0.99 1.00 1.00 1.00 0.99 0.98 1.00 0.99 0.99 0.95 0.96 0.99 1.00 0.99 0.99 0.99 0.99 0.99 0.98 0.96 0.98 1.00 0.99
Gemini 0.91 0.98 0.91 0.97 0.96 0.93 0.99 0.99 0.98 0.95 N/A 0.99 0.99 0.95 0.96 0.97 0.96 0.90 0.93 0.90 0.92 0.95 0.96
Mistral-7B-Instruct-v0.2 0.98 1.00 0.99 1.00 0.99 0.99 1.00 0.99 0.99 0.96 0.97 1.00 1.00 0.98 0.99 0.99 0.99 0.99 0.99 0.97 0.98 0.99 0.99
OPT-IML-Max-30b 0.98 0.99 0.96 0.98 0.94 0.97 0.95 0.96 0.96 0.86 0.79 0.96 0.99 0.95 0.97 0.95 0.95 0.98 0.97 0.91 0.96 0.99 0.96
Vicuna-13b 0.99 1.00 0.99 1.00 0.99 0.99 1.00 0.99 0.99 0.95 0.95 0.99 1.00 0.99 0.99 0.99 0.99 0.99 0.99 0.97 0.97 0.99 0.99
v5-Eagle-7B-HF 0.98 1.00 1.00 1.00 0.99 0.99 1.00 1.00 1.00 0.96 0.97 1.00 1.00 0.99 1.00 1.00 0.99 0.99 0.99 0.99 0.99 1.00 0.99
XLM-RoBERTa-large-MultiSocial Aya-101 0.89 0.97 0.94 0.97 0.95 0.94 0.93 0.94 0.95 0.81 0.73 0.94 0.98 0.95 0.95 0.91 0.93 0.90 0.96 0.87 0.86 0.85 0.93
GPT-3.5-Turbo-0125 0.98 0.99 0.98 0.99 0.98 0.97 0.99 0.98 0.97 0.92 0.88 0.98 0.99 0.98 0.98 0.98 0.98 0.97 0.98 0.95 0.96 0.98 0.98
Gemini 0.89 0.96 0.78 0.96 0.93 0.94 0.98 0.96 0.99 0.84 N/A 0.98 0.99 0.91 0.95 0.96 0.97 0.89 0.95 0.88 0.91 0.93 0.94
Mistral-7B-Instruct-v0.2 0.98 0.99 0.96 0.99 0.97 0.97 0.99 0.96 0.97 0.96 0.91 0.98 0.99 0.93 0.98 0.98 0.98 0.97 0.97 0.94 0.96 0.95 0.97
OPT-IML-Max-30b 0.94 0.97 0.88 0.94 0.88 0.90 0.88 0.92 0.91 0.69 0.47 0.90 0.95 0.88 0.93 0.88 0.91 0.93 0.90 0.82 0.86 0.89 0.90
Vicuna-13b 0.99 0.99 0.96 0.99 0.98 0.97 0.99 0.98 0.97 0.91 0.80 0.98 0.99 0.97 0.99 0.98 0.98 0.98 0.97 0.94 0.94 0.96 0.98
v5-Eagle-7B-HF 0.98 0.99 0.98 1.00 0.99 0.98 0.99 0.99 0.99 0.95 0.83 0.99 1.00 0.98 0.99 0.99 0.99 0.98 0.99 0.97 0.98 0.98 0.99
mDeBERTa-v3-base-MultiSocial Aya-101 0.88 0.97 0.92 0.97 0.95 0.93 0.94 0.94 0.97 0.86 0.74 0.96 0.99 0.95 0.95 0.93 0.93 0.89 0.95 0.89 0.87 0.88 0.93
GPT-3.5-Turbo-0125 0.97 0.99 0.98 0.99 0.98 0.97 0.99 0.98 0.98 0.91 0.82 0.98 0.99 0.98 0.98 0.97 0.97 0.96 0.98 0.95 0.96 0.99 0.98
Gemini 0.83 0.95 0.81 0.92 0.88 0.87 0.94 0.93 0.97 0.78 N/A 0.92 0.99 0.89 0.91 0.91 0.91 0.80 0.88 0.81 0.85 0.83 0.90
Mistral-7B-Instruct-v0.2 0.98 0.99 0.96 0.99 0.97 0.97 0.99 0.97 0.98 0.96 0.86 0.99 1.00 0.95 0.98 0.98 0.99 0.97 0.98 0.96 0.96 0.97 0.97
OPT-IML-Max-30b 0.96 0.99 0.91 0.95 0.91 0.92 0.90 0.93 0.93 0.75 0.54 0.93 0.97 0.91 0.95 0.92 0.93 0.95 0.94 0.89 0.92 0.97 0.93
Vicuna-13b 0.99 0.99 0.96 0.99 0.98 0.98 0.99 0.98 0.98 0.95 0.86 0.99 0.99 0.98 0.99 0.99 0.99 0.98 0.98 0.96 0.95 0.98 0.98
v5-Eagle-7B-HF 0.97 0.99 0.98 0.99 0.99 0.98 0.99 0.99 0.99 0.95 0.87 0.99 1.00 0.99 0.99 0.99 0.99 0.98 0.99 0.98 0.98 0.98 0.99
Table 18: Per-LLM AUC ROC performance of fine-tuned MGT detectors category. N/A refers to not enough samples per each class (at least 10) making AUC ROC value irrelevant.
Test Language [AUC ROC]
Detector Platform ar bg ca cs de el en es et ga gd hr hu nl pl pt ro ru sk sl uk zh all
Binoculars Discord N/A N/A 0.78 0.76 0.80 N/A 0.86 0.81 0.75 0.67 0.70 0.83 0.82 0.80 0.84 0.81 0.77 N/A 0.72 N/A N/A N/A 0.79
Gab 0.61 0.62 0.47 0.58 0.69 0.71 0.78 0.72 0.70 0.72 N/A 0.72 0.68 0.71 0.67 0.72 0.67 0.69 0.62 0.70 0.63 0.75 0.68
Telegram 0.71 0.69 0.61 0.77 0.74 0.83 0.81 0.77 0.75 0.72 N/A 0.79 0.82 0.72 0.78 0.76 0.79 0.65 0.72 0.77 0.63 0.73 0.73
Twitter 0.67 0.69 0.64 0.76 0.82 0.92 0.81 0.67 0.84 N/A N/A 0.85 0.79 0.72 0.82 0.80 0.77 0.85 0.87 N/A N/A 0.80 0.74
WhatsApp 0.82 N/A 0.73 0.66 0.79 N/A 0.75 0.81 0.52 N/A N/A N/A N/A 0.74 N/A 0.70 0.85 0.66 N/A N/A N/A N/A 0.73
DetectLLM-LRR Discord N/A N/A 0.95 0.98 0.91 N/A 0.92 0.93 0.94 0.83 0.75 0.94 0.98 0.94 0.98 0.95 0.96 N/A 0.90 N/A N/A N/A 0.94
Gab 0.74 0.80 0.58 0.80 0.74 0.79 0.78 0.75 0.71 0.75 N/A 0.77 0.81 0.72 0.82 0.75 0.77 0.72 0.71 0.74 0.77 0.78 0.69
Telegram 0.76 0.86 0.63 0.94 0.69 0.94 0.81 0.79 0.92 0.77 N/A 0.94 0.97 0.74 0.88 0.86 0.92 0.79 0.89 0.96 0.74 0.76 0.75
Twitter 0.81 0.87 0.75 0.91 0.91 0.96 0.85 0.78 0.87 N/A N/A 0.94 0.94 0.85 0.95 0.91 0.90 0.93 0.95 N/A N/A 0.92 0.75
WhatsApp 0.87 N/A 0.87 0.96 0.80 N/A 0.66 0.89 0.91 N/A N/A N/A N/A 0.86 N/A 0.80 0.95 0.69 N/A N/A N/A N/A 0.70
Fast-Detect-GPT Discord N/A N/A 0.64 0.81 0.81 N/A 0.86 0.82 0.70 0.76 0.76 0.81 0.79 0.81 0.83 0.84 0.80 N/A 0.65 N/A N/A N/A 0.79
Gab 0.70 0.68 0.57 0.69 0.72 0.60 0.76 0.71 0.62 0.64 N/A 0.72 0.65 0.74 0.73 0.72 0.71 0.69 0.64 0.69 0.72 0.78 0.70
Telegram 0.76 0.65 0.62 0.83 0.73 0.69 0.81 0.74 0.70 0.70 N/A 0.83 0.81 0.72 0.78 0.76 0.80 0.69 0.72 0.83 0.69 0.72 0.74
Twitter 0.71 0.65 0.62 0.82 0.82 0.78 0.81 0.65 0.67 N/A N/A 0.83 0.78 0.72 0.77 0.79 0.78 0.87 0.77 N/A N/A 0.85 0.75
WhatsApp 0.85 N/A 0.69 0.75 0.79 N/A 0.77 0.80 0.70 N/A N/A N/A N/A 0.72 N/A 0.77 0.80 0.64 N/A N/A N/A N/A 0.77
LLM-Deviation Discord N/A N/A 0.95 0.98 0.92 N/A 0.93 0.95 0.95 0.86 0.81 0.94 0.99 0.94 0.98 0.96 0.97 N/A 0.91 N/A N/A N/A 0.94
Gab 0.75 0.80 0.58 0.82 0.76 0.82 0.76 0.77 0.75 0.78 N/A 0.78 0.82 0.73 0.83 0.75 0.78 0.73 0.70 0.76 0.81 0.81 0.70
Telegram 0.79 0.85 0.62 0.94 0.72 0.93 0.81 0.79 0.92 0.80 N/A 0.95 0.98 0.76 0.89 0.86 0.93 0.79 0.90 0.97 0.74 0.77 0.74
Twitter 0.83 0.88 0.74 0.92 0.92 0.97 0.85 0.75 0.88 N/A N/A 0.94 0.95 0.82 0.96 0.92 0.91 0.93 0.96 N/A N/A 0.92 0.74
WhatsApp 0.90 N/A 0.88 0.97 0.82 N/A 0.64 0.89 0.91 N/A N/A N/A N/A 0.87 N/A 0.82 0.96 0.70 N/A N/A N/A N/A 0.70
S5 Discord N/A N/A 0.94 0.97 0.91 N/A 0.91 0.93 0.93 0.86 0.79 0.93 0.98 0.93 0.97 0.95 0.96 N/A 0.89 N/A N/A N/A 0.93
Gab 0.74 0.79 0.59 0.81 0.74 0.82 0.73 0.76 0.76 0.77 N/A 0.77 0.80 0.72 0.81 0.74 0.77 0.72 0.67 0.74 0.79 0.81 0.69
Telegram 0.78 0.84 0.62 0.93 0.70 0.93 0.79 0.77 0.91 0.77 N/A 0.94 0.97 0.75 0.88 0.85 0.92 0.79 0.89 0.95 0.74 0.76 0.73
Twitter 0.80 0.87 0.71 0.91 0.90 0.97 0.83 0.73 0.86 N/A N/A 0.94 0.93 0.78 0.95 0.91 0.90 0.93 0.94 N/A N/A 0.90 0.74
WhatsApp 0.90 N/A 0.86 0.97 0.83 N/A 0.62 0.87 0.90 N/A N/A N/A N/A 0.86 N/A 0.81 0.96 0.70 N/A N/A N/A N/A 0.70
Table 19: Per-platform AUC ROC performance of statistical MGT detectors category. N/A refers to not enough samples per each class (at least 10) making AUC ROC value irrelevant.
Test Language [AUC ROC]
Detector Platform ar bg ca cs de el en es et ga gd hr hu nl pl pt ro ru sk sl uk zh all
BLOOMZ-3b-mixed-Detector Discord N/A N/A 0.96 0.89 0.87 N/A 0.87 0.90 0.87 0.80 0.66 0.84 0.86 0.91 0.91 0.89 0.78 N/A 0.71 N/A N/A N/A 0.87
Gab 0.69 0.71 0.77 0.69 0.67 0.62 0.77 0.69 0.73 0.75 N/A 0.61 0.75 0.69 0.70 0.69 0.53 0.59 0.70 0.54 0.58 0.64 0.66
Telegram 0.81 0.79 0.73 0.81 0.76 0.85 0.83 0.80 0.84 0.78 N/A 0.80 0.88 0.74 0.77 0.80 0.75 0.75 0.76 0.84 0.71 0.71 0.78
Twitter 0.82 0.68 0.70 0.71 0.78 0.68 0.81 0.68 0.77 N/A N/A 0.81 0.71 0.70 0.70 0.80 0.59 0.73 0.90 N/A N/A 0.68 0.72
WhatsApp 0.82 N/A 0.88 0.46 0.78 N/A 0.76 0.87 0.81 N/A N/A N/A N/A 0.67 N/A 0.81 0.78 0.80 N/A N/A N/A N/A 0.79
ChatGPT-Detector-RoBERTa-Chinese Discord N/A N/A 0.88 0.72 0.86 N/A 0.95 0.88 0.85 0.73 0.62 0.66 0.82 0.71 0.70 0.75 0.83 N/A 0.68 N/A N/A N/A 0.77
Gab 0.57 0.81 0.65 0.55 0.74 0.68 0.88 0.74 0.74 0.72 N/A 0.60 0.66 0.72 0.58 0.75 0.68 0.65 0.75 0.61 0.74 0.78 0.67
Telegram 0.65 0.75 0.73 0.66 0.82 0.78 0.90 0.80 0.83 0.75 N/A 0.61 0.83 0.63 0.62 0.75 0.80 0.78 0.60 0.74 0.74 0.81 0.70
Twitter 0.86 0.90 0.73 0.67 0.77 0.78 0.90 0.79 0.72 N/A N/A 0.61 0.68 0.76 0.58 0.71 0.70 0.88 0.77 N/A N/A 0.83 0.74
WhatsApp 0.82 N/A 0.77 0.36 0.91 N/A 0.89 0.85 0.83 N/A N/A N/A N/A 0.58 N/A 0.71 0.83 0.81 N/A N/A N/A N/A 0.79
Longformer Detector Discord N/A N/A 0.25 0.41 0.47 N/A 0.68 0.44 0.40 0.50 0.46 0.50 0.40 0.42 0.44 0.36 0.33 N/A 0.36 N/A N/A N/A 0.44
Gab 0.56 0.39 0.32 0.44 0.42 0.49 0.71 0.45 0.48 0.56 N/A 0.49 0.45 0.41 0.39 0.48 0.50 0.54 0.34 0.46 0.48 0.49 0.48
Telegram 0.37 0.78 0.37 0.55 0.36 0.56 0.67 0.43 0.55 0.51 N/A 0.55 0.61 0.47 0.54 0.44 0.41 0.58 0.49 0.45 0.62 0.46 0.51
Twitter 0.24 0.16 0.30 0.38 0.49 0.67 0.65 0.31 0.43 N/A N/A 0.38 0.44 0.33 0.47 0.38 0.40 0.31 0.16 N/A N/A 0.49 0.38
WhatsApp 0.22 N/A 0.41 0.63 0.30 N/A 0.52 0.53 0.42 N/A N/A N/A N/A 0.44 N/A 0.42 0.38 0.31 N/A N/A N/A N/A 0.44
RoBERTa-large-OpenAI-Detector Discord N/A N/A 0.25 0.05 0.12 N/A 0.44 0.10 0.18 0.32 0.52 0.12 0.09 0.09 0.07 0.09 0.09 N/A 0.22 N/A N/A N/A 0.16
Gab 0.58 0.56 0.30 0.25 0.38 0.75 0.51 0.33 0.29 0.30 N/A 0.38 0.20 0.35 0.31 0.33 0.37 0.50 0.45 0.37 0.43 0.56 0.40
Telegram 0.72 0.21 0.50 0.11 0.42 0.74 0.50 0.32 0.17 0.27 N/A 0.10 0.09 0.27 0.17 0.20 0.10 0.36 0.22 0.18 0.35 0.62 0.33
Twitter 0.80 0.60 0.44 0.17 0.19 0.75 0.55 0.42 0.21 N/A N/A 0.20 0.16 0.31 0.17 0.15 0.28 0.66 0.12 N/A N/A 0.59 0.42
WhatsApp 0.82 N/A 0.26 0.20 0.32 N/A 0.55 0.17 0.21 N/A N/A N/A N/A 0.18 N/A 0.27 0.07 0.71 N/A N/A N/A N/A 0.40
ruRoBERTa-ruatd-binary Discord N/A N/A 0.51 0.41 0.25 N/A 0.53 0.42 0.35 0.41 0.36 0.45 0.43 0.43 0.40 0.35 0.25 N/A 0.32 N/A N/A N/A 0.40
Gab 0.38 0.59 0.62 0.46 0.46 0.40 0.61 0.47 0.40 0.53 N/A 0.48 0.46 0.55 0.47 0.44 0.42 0.71 0.50 0.45 0.47 0.45 0.49
Telegram 0.39 0.63 0.56 0.47 0.50 0.32 0.55 0.44 0.46 0.54 N/A 0.50 0.49 0.45 0.40 0.43 0.32 0.74 0.45 0.51 0.60 0.45 0.48
Twitter 0.39 0.67 0.55 0.31 0.38 0.23 0.58 0.49 0.45 N/A N/A 0.37 0.39 0.50 0.41 0.42 0.33 0.69 0.23 N/A N/A 0.25 0.49
WhatsApp 0.41 N/A 0.58 0.57 0.37 N/A 0.54 0.49 0.43 N/A N/A N/A N/A 0.43 N/A 0.48 0.34 0.70 N/A N/A N/A N/A 0.50
Table 20: Per-platform AUC ROC performance of pre-trained MGT detectors category. N/A refers to not enough samples per each class (at least 10) making AUC ROC value irrelevant.
Test Language [AUC ROC]
Detector Platform ar bg ca cs de el en es et ga gd hr hu nl pl pt ro ru sk sl uk zh all
Aya-101-MultiSocial Discord N/A N/A 0.99 1.00 0.97 N/A 0.99 0.99 1.00 0.96 0.90 0.99 1.00 0.99 0.99 0.99 0.99 N/A 0.99 N/A N/A N/A 0.99
Gab 0.91 0.96 0.97 0.96 0.96 0.95 0.96 0.96 0.94 0.94 N/A 0.94 0.98 0.97 0.97 0.96 0.95 0.92 0.96 0.92 0.89 0.94 0.94
Telegram 0.97 0.99 0.97 0.98 0.98 0.98 0.98 0.97 0.99 0.94 N/A 0.99 1.00 0.96 0.98 0.97 0.99 0.97 0.98 1.00 0.96 0.99 0.98
Twitter 0.99 0.99 0.98 0.97 0.97 0.99 0.97 0.99 0.98 N/A N/A 1.00 0.98 0.97 0.99 0.98 0.98 0.98 1.00 N/A N/A 0.97 0.98
WhatsApp 0.99 N/A 0.97 0.98 0.95 N/A 0.98 0.98 1.00 N/A N/A N/A N/A 0.97 N/A 0.97 0.99 0.95 N/A N/A N/A N/A 0.97
BLOOMZ-3b-MultiSocial Discord N/A N/A 0.99 1.00 0.97 N/A 0.99 0.99 0.99 0.97 0.84 0.98 1.00 0.98 0.99 0.99 0.99 N/A 0.97 N/A N/A N/A 0.99
Gab 0.87 0.96 0.96 0.93 0.92 0.93 0.96 0.95 0.94 0.88 N/A 0.90 0.97 0.92 0.94 0.95 0.89 0.88 0.95 0.81 0.81 0.93 0.91
Telegram 0.95 0.98 0.94 0.97 0.96 0.98 0.98 0.97 0.98 0.86 N/A 0.97 1.00 0.92 0.94 0.97 0.97 0.94 0.95 0.98 0.90 0.98 0.96
Twitter 0.99 0.98 0.98 0.94 0.96 0.99 0.98 0.99 0.92 N/A N/A 0.99 0.97 0.94 0.97 0.97 0.96 0.99 0.98 N/A N/A 0.99 0.97
WhatsApp 0.99 N/A 0.97 0.98 0.97 N/A 0.98 0.98 0.99 N/A N/A N/A N/A 0.95 N/A 0.98 0.99 0.97 N/A N/A N/A N/A 0.98
Falcon-rw-1b-MultiSocial Discord N/A N/A 0.99 1.00 0.97 N/A 0.99 0.98 0.99 0.96 0.87 0.98 1.00 0.99 0.99 0.99 0.99 N/A 0.97 N/A N/A N/A 0.99
Gab 0.88 0.94 0.97 0.94 0.93 0.94 0.97 0.93 0.94 0.92 N/A 0.91 0.97 0.93 0.95 0.92 0.89 0.88 0.93 0.80 0.87 0.92 0.92
Telegram 0.94 0.98 0.93 0.97 0.96 0.98 0.98 0.96 0.99 0.89 N/A 0.98 1.00 0.94 0.95 0.95 0.98 0.94 0.96 0.98 0.91 0.97 0.96
Twitter 0.98 0.98 0.98 0.95 0.96 0.99 0.98 0.98 0.95 N/A N/A 0.97 0.96 0.95 0.96 0.97 0.96 0.99 0.99 N/A N/A 0.99 0.97
WhatsApp 0.99 N/A 0.95 0.97 0.96 N/A 0.99 0.97 0.99 N/A N/A N/A N/A 0.95 N/A 0.96 0.99 0.96 N/A N/A N/A N/A 0.97
Llama-3-8b-MultiSocial Discord N/A N/A 0.99 1.00 0.97 N/A 0.99 0.99 1.00 0.96 0.88 0.99 1.00 0.99 1.00 0.99 0.99 N/A 0.98 N/A N/A N/A 0.99
Gab 0.90 0.95 0.98 0.97 0.97 0.95 0.98 0.97 0.95 0.94 N/A 0.96 0.99 0.97 0.97 0.96 0.95 0.93 0.97 0.92 0.88 0.96 0.95
Telegram 0.97 0.99 0.97 0.99 0.98 0.98 0.98 0.98 0.99 0.94 N/A 0.99 1.00 0.97 0.98 0.97 0.98 0.97 0.98 1.00 0.96 0.99 0.98
Twitter 0.99 0.99 0.99 0.97 0.98 0.98 0.99 0.99 0.97 N/A N/A 0.99 0.98 0.98 0.99 0.98 0.98 0.98 1.00 N/A N/A 0.98 0.98
WhatsApp 0.99 N/A 0.98 0.98 0.96 N/A 0.99 0.98 1.00 N/A N/A N/A N/A 0.98 N/A 0.98 0.99 0.98 N/A N/A N/A N/A 0.98
Mistral-7b-v0.1-MultiSocial Discord N/A N/A 0.99 1.00 0.97 N/A 0.99 0.99 1.00 0.96 0.91 0.99 1.00 0.99 1.00 0.99 0.99 N/A 0.98 N/A N/A N/A 0.99
Gab 0.90 0.97 0.97 0.98 0.97 0.96 0.98 0.97 0.94 0.92 N/A 0.96 0.99 0.97 0.98 0.96 0.94 0.94 0.95 0.91 0.89 0.96 0.95
Telegram 0.97 0.99 0.97 0.99 0.98 0.98 0.98 0.98 0.99 0.91 N/A 0.99 1.00 0.97 0.98 0.97 0.99 0.97 0.97 1.00 0.96 0.99 0.98
Twitter 0.99 0.99 0.98 0.98 0.98 0.99 0.99 0.99 0.96 N/A N/A 0.99 0.98 0.97 0.98 0.98 0.98 0.99 0.99 N/A N/A 0.98 0.98
WhatsApp 0.99 N/A 0.97 0.98 0.96 N/A 0.99 0.98 0.99 N/A N/A N/A N/A 0.97 N/A 0.98 0.99 0.98 N/A N/A N/A N/A 0.98
XLM-RoBERTa-large-MultiSocial Discord N/A N/A 0.96 0.99 0.96 N/A 0.98 0.97 0.99 0.89 0.72 0.98 1.00 0.98 0.99 0.97 0.98 N/A 0.99 N/A N/A N/A 0.97
Gab 0.88 0.94 0.95 0.93 0.94 0.94 0.95 0.95 0.90 0.87 N/A 0.94 0.96 0.95 0.95 0.93 0.94 0.90 0.95 0.88 0.84 0.92 0.93
Telegram 0.94 0.98 0.93 0.98 0.97 0.97 0.96 0.96 0.98 0.87 N/A 0.98 0.99 0.94 0.97 0.96 0.97 0.96 0.96 0.98 0.93 0.94 0.96
Twitter 0.98 0.99 0.94 0.96 0.96 0.97 0.97 0.97 0.97 N/A N/A 0.95 0.97 0.94 0.98 0.96 0.97 0.98 0.98 N/A N/A 0.96 0.96
WhatsApp 0.98 N/A 0.90 0.96 0.97 N/A 0.97 0.97 0.97 N/A N/A N/A N/A 0.93 N/A 0.95 0.97 0.95 N/A N/A N/A N/A 0.96
mDeBERTa-v3-base-MultiSocial Discord N/A N/A 0.98 1.00 0.97 N/A 0.98 0.98 0.99 0.91 0.75 0.98 1.00 0.98 0.99 0.99 0.99 N/A 0.97 N/A N/A N/A 0.98
Gab 0.85 0.96 0.92 0.94 0.93 0.90 0.95 0.93 0.92 0.88 N/A 0.92 0.97 0.94 0.94 0.92 0.92 0.89 0.92 0.88 0.79 0.91 0.91
Telegram 0.94 0.99 0.93 0.98 0.95 0.97 0.97 0.95 0.98 0.90 N/A 0.98 1.00 0.93 0.96 0.96 0.97 0.96 0.96 0.99 0.94 0.96 0.96
Twitter 0.98 0.98 0.96 0.93 0.95 0.97 0.96 0.97 0.97 N/A N/A 0.96 0.96 0.95 0.97 0.95 0.96 0.96 0.99 N/A N/A 0.95 0.96
WhatsApp 0.97 N/A 0.94 0.96 0.96 N/A 0.96 0.97 1.00 N/A N/A N/A N/A 0.96 N/A 0.95 0.99 0.95 N/A N/A N/A N/A 0.96
Table 21: Per-platform AUC ROC performance of fine-tuned MGT detectors category. N/A refers to not enough samples per each class (at least 10) making AUC ROC value irrelevant.
Platform [AUC ROC]
Detector Generator Discord Gab Telegram Twitter WhatsApp all
Binoculars Mistral-7B-Instruct-v0.2 0.77 0.62 0.68 0.68 0.67 0.68
aya-101 0.76 0.65 0.68 0.70 0.72 0.69
gemini 0.88 0.79 0.84 0.83 0.84 0.83
gpt-3.5-turbo-0125 0.75 0.64 0.71 0.68 0.69 0.68
opt-iml-max-30b 0.71 0.60 0.64 0.66 0.65 0.64
v5-Eagle-7B-HF 0.85 0.76 0.80 0.81 0.79 0.79
vicuna-13b 0.82 0.73 0.76 0.80 0.77 0.76
DetectLLM-LRR Mistral-7B-Instruct-v0.2 0.94 0.63 0.71 0.71 0.68 0.71
aya-101 0.91 0.64 0.69 0.70 0.66 0.70
gemini 0.96 0.80 0.84 0.82 0.78 0.83
gpt-3.5-turbo-0125 0.93 0.65 0.72 0.71 0.67 0.71
opt-iml-max-30b 0.87 0.60 0.68 0.67 0.63 0.67
v5-Eagle-7B-HF 0.98 0.78 0.81 0.83 0.77 0.82
vicuna-13b 0.96 0.73 0.77 0.80 0.73 0.78
Fast-Detect-GPT Mistral-7B-Instruct-v0.2 0.65 0.51 0.58 0.59 0.62 0.58
aya-101 0.80 0.71 0.74 0.75 0.77 0.75
gemini 0.89 0.84 0.88 0.86 0.88 0.87
gpt-3.5-turbo-0125 0.75 0.66 0.73 0.68 0.74 0.71
opt-iml-max-30b 0.75 0.61 0.65 0.66 0.71 0.67
v5-Eagle-7B-HF 0.88 0.81 0.84 0.86 0.86 0.85
vicuna-13b 0.81 0.74 0.75 0.82 0.82 0.78
LLM-Deviation Mistral-7B-Instruct-v0.2 0.95 0.63 0.69 0.70 0.67 0.70
aya-101 0.92 0.65 0.70 0.70 0.66 0.70
gemini 0.97 0.80 0.82 0.82 0.78 0.82
gpt-3.5-turbo-0125 0.93 0.65 0.71 0.69 0.67 0.71
opt-iml-max-30b 0.89 0.62 0.69 0.67 0.64 0.68
v5-Eagle-7B-HF 0.98 0.78 0.80 0.83 0.77 0.82
vicuna-13b 0.96 0.73 0.76 0.80 0.73 0.78
S5 Mistral-7B-Instruct-v0.2 0.93 0.63 0.69 0.69 0.67 0.69
aya-101 0.91 0.65 0.70 0.70 0.66 0.71
gemini 0.96 0.78 0.81 0.81 0.77 0.81
gpt-3.5-turbo-0125 0.91 0.64 0.70 0.68 0.66 0.70
opt-iml-max-30b 0.87 0.62 0.69 0.67 0.64 0.69
v5-Eagle-7B-HF 0.98 0.78 0.80 0.83 0.77 0.81
vicuna-13b 0.96 0.73 0.75 0.80 0.73 0.78
Table 22: Per-platform per-LLM AUC ROC performance of statistical MGT detectors category. N/A refers to not enough samples per each class (at least 10) making AUC ROC value irrelevant.
Platform [AUC ROC]
Detector Generator Discord Gab Telegram Twitter WhatsApp all
BLOOMZ-3b-mixed-Detector Mistral-7B-Instruct-v0.2 0.87 0.62 0.76 0.70 0.78 0.74
aya-101 0.92 0.75 0.85 0.81 0.85 0.83
gemini 0.76 0.47 0.63 0.52 0.58 0.59
gpt-3.5-turbo-0125 0.91 0.70 0.80 0.75 0.86 0.80
opt-iml-max-30b 0.81 0.67 0.76 0.69 0.79 0.73
v5-Eagle-7B-HF 0.92 0.70 0.83 0.78 0.86 0.81
vicuna-13b 0.89 0.68 0.81 0.78 0.83 0.79
ChatGPT-Detector-RoBERTa-Chinese Mistral-7B-Instruct-v0.2 0.85 0.70 0.75 0.74 0.85 0.76
aya-101 0.70 0.64 0.65 0.71 0.73 0.67
gemini 0.88 0.73 0.78 0.81 0.86 0.80
gpt-3.5-turbo-0125 0.72 0.59 0.66 0.65 0.73 0.66
opt-iml-max-30b 0.65 0.63 0.64 0.70 0.72 0.65
v5-Eagle-7B-HF 0.80 0.71 0.74 0.80 0.82 0.76
vicuna-13b 0.76 0.68 0.71 0.77 0.81 0.72
Longformer Detector Mistral-7B-Instruct-v0.2 0.51 0.55 0.59 0.44 0.47 0.52
aya-101 0.32 0.38 0.40 0.28 0.37 0.36
gemini 0.50 0.54 0.58 0.47 0.52 0.53
gpt-3.5-turbo-0125 0.37 0.50 0.49 0.44 0.44 0.45
opt-iml-max-30b 0.36 0.38 0.40 0.27 0.34 0.36
v5-Eagle-7B-HF 0.53 0.55 0.60 0.41 0.52 0.53
vicuna-13b 0.43 0.48 0.53 0.38 0.44 0.46
RoBERTa-large-OpenAI-Detector Mistral-7B-Instruct-v0.2 0.11 0.38 0.30 0.38 0.33 0.31
aya-101 0.25 0.48 0.40 0.49 0.52 0.43
gemini 0.07 0.25 0.22 0.29 0.21 0.21
gpt-3.5-turbo-0125 0.17 0.40 0.33 0.42 0.38 0.35
opt-iml-max-30b 0.29 0.56 0.47 0.57 0.57 0.49
v5-Eagle-7B-HF 0.11 0.36 0.30 0.38 0.40 0.31
vicuna-13b 0.13 0.39 0.33 0.40 0.41 0.33
ruRoBERTa-ruatd-binary Mistral-7B-Instruct-v0.2 0.47 0.52 0.54 0.53 0.53 0.53
aya-101 0.42 0.57 0.52 0.53 0.57 0.53
gemini 0.16 0.17 0.16 0.20 0.19 0.18
gpt-3.5-turbo-0125 0.40 0.52 0.49 0.53 0.56 0.50
opt-iml-max-30b 0.45 0.63 0.60 0.62 0.64 0.59
v5-Eagle-7B-HF 0.46 0.50 0.53 0.49 0.53 0.52
vicuna-13b 0.41 0.51 0.53 0.51 0.53 0.51
Table 23: Per-platform per-LLM AUC ROC performance of pre-trained MGT detectors category. N/A refers to not enough samples per each class (at least 10) making AUC ROC value irrelevant.
Platform [AUC ROC]
Detector Generator Discord Gab Telegram Twitter WhatsApp all
Aya-101-MultiSocial Mistral-7B-Instruct-v0.2 1.00 0.98 0.99 0.99 1.00 0.99
aya-101 0.99 0.91 0.97 0.96 0.95 0.96
gemini 0.98 0.86 0.95 0.93 0.93 0.93
gpt-3.5-turbo-0125 1.00 0.98 0.99 1.00 1.00 0.99
opt-iml-max-30b 0.97 0.92 0.97 0.97 0.96 0.95
v5-Eagle-7B-HF 1.00 0.98 1.00 1.00 0.99 0.99
vicuna-13b 1.00 0.98 0.99 1.00 1.00 0.99
BLOOMZ-3b-MultiSocial Mistral-7B-Instruct-v0.2 1.00 0.94 0.97 0.98 0.99 0.97
aya-101 0.98 0.88 0.94 0.95 0.96 0.94
gemini 0.99 0.85 0.94 0.93 0.96 0.93
gpt-3.5-turbo-0125 0.99 0.95 0.98 0.99 0.99 0.98
opt-iml-max-30b 0.96 0.87 0.94 0.95 0.95 0.93
v5-Eagle-7B-HF 1.00 0.96 0.99 0.99 0.99 0.99
vicuna-13b 0.99 0.95 0.98 0.99 0.99 0.98
Falcon-rw-1b-MultiSocial Mistral-7B-Instruct-v0.2 1.00 0.93 0.97 0.98 0.98 0.97
aya-101 0.98 0.89 0.95 0.96 0.96 0.94
gemini 0.98 0.84 0.94 0.94 0.94 0.93
gpt-3.5-turbo-0125 0.99 0.96 0.98 0.99 0.99 0.98
opt-iml-max-30b 0.96 0.88 0.94 0.95 0.95 0.93
v5-Eagle-7B-HF 1.00 0.96 0.99 0.99 0.99 0.98
vicuna-13b 0.99 0.95 0.97 0.99 0.99 0.98
Llama-3-8b-MultiSocial Mistral-7B-Instruct-v0.2 1.00 0.98 0.99 0.99 1.00 0.99
aya-101 0.99 0.92 0.97 0.97 0.97 0.96
gemini 0.99 0.89 0.95 0.94 0.95 0.95
gpt-3.5-turbo-0125 1.00 0.98 0.99 1.00 1.00 0.99
opt-iml-max-30b 0.97 0.93 0.97 0.97 0.97 0.96
v5-Eagle-7B-HF 1.00 0.99 1.00 1.00 1.00 0.99
vicuna-13b 1.00 0.98 0.99 1.00 1.00 0.99
Mistral-7b-v0.1-MultiSocial Mistral-7B-Instruct-v0.2 1.00 0.97 0.99 0.99 1.00 0.99
aya-101 0.99 0.92 0.97 0.96 0.97 0.96
gemini 0.99 0.91 0.96 0.95 0.97 0.96
gpt-3.5-turbo-0125 1.00 0.98 0.99 1.00 1.00 0.99
opt-iml-max-30b 0.97 0.92 0.97 0.97 0.97 0.96
v5-Eagle-7B-HF 1.00 0.98 1.00 1.00 1.00 0.99
vicuna-13b 1.00 0.98 0.99 0.99 1.00 0.99
XLM-RoBERTa-large-MultiSocial Mistral-7B-Instruct-v0.2 0.99 0.95 0.97 0.97 0.99 0.97
aya-101 0.96 0.89 0.93 0.95 0.92 0.93
gemini 0.97 0.89 0.96 0.94 0.96 0.94
gpt-3.5-turbo-0125 0.98 0.96 0.98 0.99 0.99 0.98
opt-iml-max-30b 0.91 0.87 0.91 0.93 0.91 0.90
v5-Eagle-7B-HF 1.00 0.97 0.99 0.99 0.99 0.99
vicuna-13b 0.99 0.96 0.97 0.99 0.99 0.98
mDeBERTa-v3-base-MultiSocial Mistral-7B-Instruct-v0.2 0.99 0.94 0.98 0.98 0.99 0.97
aya-101 0.98 0.88 0.94 0.94 0.93 0.93
gemini 0.97 0.80 0.92 0.88 0.90 0.90
gpt-3.5-turbo-0125 0.99 0.95 0.98 0.98 0.99 0.98
opt-iml-max-30b 0.95 0.89 0.94 0.94 0.94 0.93
v5-Eagle-7B-HF 1.00 0.97 0.99 0.99 0.99 0.99
vicuna-13b 0.99 0.96 0.98 0.99 0.99 0.98
Table 24: Per-platform per-LLM AUC ROC performance of fine-tuned MGT detectors category. N/A refers to not enough samples per each class (at least 10) making AUC ROC value irrelevant.
Train Test Language [AUC ROC]
Detector Language ar bg ca cs de el en es et ga gd hr hu nl pl pt ro ru sk sl uk zh all
Aya-101 en 0.89 0.97 0.84 0.97 0.92 0.95 0.97 0.93 0.96 N/A N/A 0.96 0.99 0.87 0.95 0.95 0.97 0.93 N/A N/A 0.90 0.91 0.93
es 0.91 0.97 0.90 0.96 0.95 0.93 0.95 0.96 0.96 N/A N/A 0.96 0.99 0.90 0.95 0.95 0.97 0.95 N/A N/A 0.90 0.91 0.94
ru 0.92 0.98 0.89 0.95 0.92 0.94 0.95 0.93 0.96 N/A N/A 0.96 0.98 0.89 0.95 0.93 0.97 0.97 N/A N/A 0.94 0.91 0.94
en-es-ru 0.93 0.98 0.90 0.96 0.94 0.94 0.96 0.95 0.96 N/A N/A 0.96 0.99 0.89 0.95 0.95 0.97 0.96 N/A N/A 0.91 0.92 0.94
BLOOMZ-3b en 0.75 0.91 0.73 0.89 0.87 0.90 0.95 0.74 0.88 N/A N/A 0.90 0.95 0.84 0.87 0.90 0.92 0.87 N/A N/A 0.82 0.61 0.82
es 0.78 0.85 0.84 0.84 0.85 0.82 0.90 0.93 0.83 N/A N/A 0.82 0.88 0.80 0.80 0.90 0.88 0.85 N/A N/A 0.79 0.61 0.81
ru 0.69 0.86 0.57 0.80 0.80 0.76 0.81 0.56 0.80 N/A N/A 0.78 0.83 0.73 0.75 0.74 0.74 0.90 N/A N/A 0.84 0.64 0.72
en-es-ru 0.87 0.90 0.83 0.87 0.89 0.92 0.95 0.92 0.86 N/A N/A 0.85 0.92 0.79 0.85 0.92 0.90 0.90 N/A N/A 0.86 0.88 0.86
Falcon-rw-1b en 0.74 0.74 0.78 0.85 0.86 0.87 0.95 0.87 0.88 N/A N/A 0.91 0.95 0.79 0.85 0.89 0.92 0.81 N/A N/A 0.74 0.80 0.83
es 0.78 0.73 0.79 0.85 0.89 0.91 0.86 0.93 0.86 N/A N/A 0.90 0.92 0.82 0.87 0.92 0.93 0.85 N/A N/A 0.75 0.80 0.83
ru 0.68 0.85 0.71 0.76 0.78 0.86 0.82 0.83 0.82 N/A N/A 0.84 0.80 0.71 0.82 0.80 0.88 0.90 N/A N/A 0.84 0.75 0.78
en-es-ru 0.82 0.87 0.83 0.91 0.90 0.93 0.96 0.92 0.91 N/A N/A 0.93 0.97 0.84 0.89 0.92 0.94 0.89 N/A N/A 0.84 0.87 0.89
Llama-3-8b en 0.85 0.96 0.76 0.89 0.89 0.90 0.97 0.91 0.93 N/A N/A 0.96 0.98 0.84 0.92 0.94 0.93 0.91 N/A N/A 0.85 0.73 0.87
es 0.80 0.94 0.85 0.79 0.90 0.88 0.91 0.95 0.86 N/A N/A 0.90 0.95 0.84 0.92 0.92 0.92 0.90 N/A N/A 0.86 0.64 0.83
ru 0.76 0.95 0.67 0.79 0.84 0.88 0.86 0.81 0.76 N/A N/A 0.83 0.89 0.80 0.90 0.83 0.82 0.95 N/A N/A 0.90 0.59 0.78
en-es-ru 0.92 0.98 0.89 0.93 0.95 0.95 0.97 0.96 0.95 N/A N/A 0.97 0.99 0.89 0.95 0.95 0.95 0.96 N/A N/A 0.93 0.93 0.93
Mistral-7b-v0.1 en 0.77 0.83 0.81 0.86 0.89 0.85 0.96 0.86 0.89 N/A N/A 0.92 0.92 0.82 0.81 0.88 0.92 0.80 N/A N/A 0.75 0.46 0.82
es 0.77 0.82 0.86 0.85 0.89 0.75 0.83 0.93 0.89 N/A N/A 0.91 0.94 0.80 0.78 0.91 0.91 0.82 N/A N/A 0.76 0.51 0.82
ru 0.78 0.94 0.76 0.89 0.82 0.85 0.82 0.84 0.89 N/A N/A 0.92 0.90 0.74 0.80 0.87 0.88 0.95 N/A N/A 0.90 0.44 0.82
en-es-ru 0.90 0.94 0.87 0.93 0.91 0.92 0.95 0.93 0.93 N/A N/A 0.95 0.97 0.85 0.86 0.93 0.94 0.96 N/A N/A 0.91 0.68 0.90
XLM-RoBERTa-large en 0.82 0.92 0.77 0.94 0.88 0.92 0.96 0.91 0.95 N/A N/A 0.96 0.99 0.87 0.91 0.94 0.96 0.88 N/A N/A 0.81 0.83 0.90
es 0.87 0.95 0.86 0.95 0.91 0.94 0.92 0.94 0.95 N/A N/A 0.97 0.99 0.86 0.91 0.94 0.97 0.90 N/A N/A 0.83 0.85 0.91
ru 0.92 0.98 0.87 0.96 0.89 0.95 0.92 0.91 0.96 N/A N/A 0.97 0.98 0.86 0.94 0.93 0.96 0.96 N/A N/A 0.92 0.91 0.93
en-es-ru 0.91 0.97 0.88 0.97 0.92 0.96 0.96 0.93 0.97 N/A N/A 0.98 0.99 0.90 0.94 0.95 0.97 0.95 N/A N/A 0.90 0.90 0.94
mDeBERTa-v3-base en 0.83 0.96 0.76 0.97 0.84 0.93 0.96 0.88 0.94 N/A N/A 0.96 0.99 0.84 0.92 0.94 0.96 0.90 N/A N/A 0.84 0.74 0.90
es 0.86 0.95 0.82 0.96 0.87 0.87 0.94 0.92 0.94 N/A N/A 0.95 0.98 0.85 0.93 0.94 0.95 0.91 N/A N/A 0.87 0.82 0.90
ru 0.89 0.96 0.82 0.96 0.85 0.92 0.94 0.88 0.95 N/A N/A 0.94 0.99 0.84 0.95 0.93 0.94 0.95 N/A N/A 0.91 0.87 0.91
en-es-ru 0.88 0.96 0.82 0.96 0.86 0.90 0.95 0.90 0.95 N/A N/A 0.95 0.99 0.86 0.93 0.94 0.95 0.93 N/A N/A 0.89 0.85 0.91
Table 25: Cross-lingual AUC ROC performance of the selected MGT detectors fine-tuned monolingually (en, es and ru) and multilingually (en-es-ru), evaluated based on Telegram data (for training as well as for testing). N/A refers to not enough samples (at least 2000) in MultiSocial Telegram data.
Train Test Platform [AUC ROC]
Detector Platform Discord Gab Telegram Twitter WhatsApp all
Aya-101 Discord 0.99 0.82 0.89 0.81 0.92 0.89
Gab 0.97 0.96 0.95 0.96 0.94 0.96
Telegram 0.98 0.92 0.97 0.95 0.96 0.95
Twitter 0.98 0.92 0.94 0.98 0.94 0.95
WhatsApp 0.97 0.88 0.94 0.93 0.98 0.94
all 0.98 0.95 0.97 0.97 0.97 0.97
BLOOMZ-3b Discord 0.98 0.85 0.87 0.84 0.88 0.88
Gab 0.93 0.92 0.89 0.91 0.89 0.91
Telegram 0.97 0.90 0.94 0.92 0.93 0.93
Twitter 0.95 0.89 0.88 0.98 0.87 0.91
WhatsApp 0.96 0.90 0.91 0.92 0.97 0.93
all 0.98 0.94 0.95 0.97 0.96 0.96
Falcon-rw-1b Discord 0.98 0.80 0.87 0.82 0.85 0.86
Gab 0.95 0.93 0.92 0.95 0.91 0.93
Telegram 0.98 0.90 0.95 0.95 0.93 0.94
Twitter 0.97 0.91 0.93 0.98 0.92 0.94
WhatsApp 0.97 0.91 0.93 0.95 0.98 0.94
all 0.97 0.91 0.94 0.96 0.94 0.94
Llama-3-8b Discord 0.99 0.83 0.82 0.72 0.86 0.83
Gab 0.95 0.96 0.93 0.97 0.90 0.94
Telegram 0.98 0.94 0.97 0.97 0.96 0.97
Twitter 0.96 0.90 0.92 0.98 0.91 0.93
WhatsApp 0.97 0.89 0.93 0.93 0.98 0.94
all 0.98 0.94 0.95 0.97 0.95 0.96
Mistral-7b-v0.1 Discord 0.98 0.82 0.88 0.86 0.88 0.88
Gab 0.95 0.95 0.93 0.96 0.92 0.94
Telegram 0.98 0.94 0.97 0.97 0.95 0.96
Twitter 0.96 0.90 0.90 0.98 0.91 0.93
WhatsApp 0.96 0.90 0.92 0.94 0.97 0.93
all 0.96 0.92 0.93 0.96 0.94 0.94
XLM-RoBERTa-large Discord 0.99 0.87 0.90 0.83 0.94 0.90
Gab 0.97 0.93 0.92 0.90 0.90 0.92
Telegram 0.98 0.92 0.96 0.93 0.95 0.95
Twitter 0.98 0.93 0.94 0.97 0.94 0.95
WhatsApp 0.97 0.88 0.92 0.88 0.96 0.92
all 0.98 0.94 0.96 0.96 0.96 0.96
mDeBERTa-v3-base Discord 0.99 0.86 0.89 0.84 0.93 0.89
Gab 0.97 0.93 0.93 0.92 0.94 0.94
Telegram 0.98 0.90 0.95 0.93 0.95 0.94
Twitter 0.97 0.90 0.92 0.97 0.94 0.94
WhatsApp 0.98 0.91 0.93 0.90 0.97 0.93
all 0.98 0.91 0.94 0.95 0.95 0.95
Table 26: Cross-platform evaluation of the selected fine-tuned MGT detectors.